### v1

Abstract
Geometric shape classification of vector polygons remains a non-trivial learning task in spatial analysis. Previous studies mainly focus on devising deep learning approaches for representation learning of rasterized vector polygons, whereas the study of discrete representations of polygons and subsequent deep learning approaches have not been fully investigated. In this study, we investigate a graph representation of vector polygons and propose a novel graph message-passing neural network (PolyMP) to learn the geometric-invariant features for shape classification of polygons. Through extensive experiments, we show that the graph representation of polygons combined with a permutation-invariant graph message-passing neural network achieves highly robust performances on benchmark datasets (i.e., synthetic glyph and real-world building footprint datasets) as compared to baseline methods. We demonstrate that the proposed graph-based PolyMP network enables the learning of expressive geometric features invariant to geometric transformations of polygons (i.e., translation, rotation, scaling and shearing) and is robust to trivial vertex removals of polygons. We further show the strong generalizability of PolyMP, which enables generalizing the learned geometric features from the synthetic glyph polygons to the real-world building footprints.
keywords:
Spatial Vector Polygons; Geometric Shape Classification; Transformation Invariance; Permutation Invariance; Deep Neural Networks
1 Introduction
Geometric shape classification of spatial objects is a non-trivial task in spatial analysis. The recognition of spatial objects assisted by automated shape classification is a major enabler of data intensive tasks, including cartographic generalisation, building pattern recognition, archaeological feature analysis, and road geometry identification [1, 2, 3].
One of the main challenges in geometric shape classification is the identification of object footprints (i.e., outlines) in the geographic context. A key requirement is the classification of objects invariant to geometric transformations including rotation, scaling, and shearing. The human visual system relies on the Gestalt properties of perceived objects [4, 5] during the classification of shapes. A key property of Gestalt is the invariance to geometric transformations, assuring that geometric shapes are recognized regardless of geometric transformations. In contrast to the human visual perception capabilities, deep learning architectures have been designed and shown to be mainly translation-invariant (e.g., Convolution Neural Networks, CNN) or permutation-invariant (e.g., Graph Neural Networks, GNN) on classification tasks. The ability to reflect Gestalt principles through transformation invariances thus present a strong motivation for an inductive bias in the design of learning-based models for geometric shape recognition of spatial objects.
Spatial objects are often conveniently represented as vector polygons, a data representation thus far neglected in deep learning research. Learning geometric-invariant features from vector polygons has the following requirements: (1) a generic data representation that encodes geometric features of polygons without information loss; and (2) a learning model built on this input data representation that enables learning latent geometric-invariant features robust to geometric transformations. Existing geospatial applications (i.e., shape coding and retrieval [6], building patter recognition [2, 1] and building grouping [7]) utilizing polygonal geometries motivate us to systemically study geometry encoding methods in conjunction with appropriate and robust learning architectures that enable learning transformation invariant features of spatial polygon geometries.
Veer et al. [2] proposed VeerCNN, a deep convolution model to learn convolutional features on fixed-size 1D vertex sequences of polygon vertices for building attribute recognition. The architectural limitations of CNNs attributable to shared-weights convolutional kernels followed by non-linear activation and fixed-size pooling layers (e.g, $mean$, $sum$ or $max$) only enable learning intermediate hierarchical features invariant to translation, with poor handling of rotation, scaling and shearing. Mai et al. [8] noted that CNN models learning on 1D sequences are also sensitive to (1) permutations of the feed-in order of polygon vertices (i.e., sensitive to loop origin invariance); and (2) the impact of trivial vertices on the exterior and interiors of polygons, defined by Mai et al. [8] as “[vertices] …where the addition or removal of the vertex have no effect on the geometric shape and topological properties of the outlines of the polygons”. In this study, we treat the addition or removal of trivial vertices of polygons as changes that do not alter the semantic information of the shape (captured by the human-assigned label), thus closely reflecting the relationship with the Gestalt principles.
Learning models tailored for discrete (i.e., not grid-like), permutation invariant data representations are predicated on data domains where entries hold no explicit neighborhood information (e.g., point sets [9, 10, 11, 12]). Polygons can then be encoded into point sets of vertices on the exterior and interior rings of the polygons with arbitrary ordering. In contrast to the 1D sequence encoding of polygons, point set encoding assumes input data to have varying input sizes and feed-in order. Models learning on point sets are designed to map input points with permuted order to task-specific outputs (i.e., permutation invariance [13]). Here, we argue that such point set representations do not sufficiently capture the connectivity information between vertices of polygons, leading to information loss and performance degradation in geometric shape classification.
Graph data structures present a suitable data encoding for polygons. Recent studies [6, 14, 15] convert vector polygons into un-directed graphs where the vertices on the exteriors and interiors of polygons are captured as graph nodes linked by un-directed graph edges. Compared to fixed-size 1D sequences and varying-size point-set encoding of polygons, graph encoding effectively encapsulates both the geometric and connectivity information of the vertices along the exterior and interior linear rings defining the polygons. Graph-based representations also enable the feed-in order of polygon vertices to be independent of the model outcomes, while the connectivity (topology) between parts of polygon representations remains invariant to geometric transformations.
Graph convolutional autoencoders (GCAE) [6] extend graph convolution neural networks [16] by learning spectral convolution embeddings from polygon graphs and demonstrate the effectiveness of latent embedding. Bronstein et al. [13] discussed that the graph convolution layers aggregate node features from neighboring nodes with constant weights, largely limited by the topological structure of the input graphs. In vector polygons vertices on the linear rings have constant neighbours (i.e., left and right neighbour vertices on linear rings), leading to a reduction in the expressivity of graph convolution features learned from polygon graphs. In this study, we propose to leverage graph message-passing mechanisms [17] in graph-based learning models to learn highly expressive and robust latent features of polygons.
Our hypothesis is that graph representations of polygons enable the learning of robust latent features of polygons invariant to geometric transformations (i.e., rotation, scaling and shearing) when combined with graph message-passing models, thus supporting generalizability on data with domain shifts.
If the hypothesis holds, we suggest that the downstream tasks performed on the learned features of vector polygons lead to robust performance on shapes with varying amounts of trivial vertices. We present the results of a series of experiments designed to evaluate model robustness on polygons altered by geometric transformations, and investigate the generalizability of the findings on combinations of polygon representations and model architectures (permutation-invariant vs. translation-invariant models) for the classification of vector polygons with and without holes (inner linear rings).
After a literature review of related literature (Section 2) we present a synthetic dataset of polygonal shapes of high shape variability with known, robust human labels, based on latin alphabet character glyphs (Section 3.1.1). Despite a broad range of shape variations between fonts (in particular with respect to trivial vertices), characters must maintain strong recognisability attributable to the shapes’ Gestalt. We then experimentally benchmark the performance of learning models operating on three distinct data representations (1D sequence, set and graph) and varying amounts of geometric transformations (i.e., rotation, scaling and shearing). Our results (Section 4) document the models’ robustness to geometric transformations by benchmarking their performance on the synthetic glyph dataset. Building on these results, we further evaluate the generalizability of the models on a real-world building footprint dataset collected from the Open Street Map (OSM) [6].
Our main contributions are:
  1. 1.
    a thorough investigation of the impact of distinct combinations of graph representation of polygons and message-passing learning architectures for the learning of geometric features robust to geometric transformations of shapes and removals of trivial vertices, grounded in a theoretically grounded set of desiderata;
  2. 2.
    the contribution of an open synthetic dataset of vector polygon shapes based on glyphs, enabling the benchmarking of geometric learning models; and
  3. 3.
    the experimental demonstration that graph representations of vector polygons in conjunction with message-passing neural networks achieve the most robust performance and generalization to real-world shape datasets.
2 Background
2.1 Machine Learning with Vector Geometries
In a straightforward extension of computer vision approaches devised for image classification, current research on polygonal shape classification typically learns salient geometric features of shapes based on rasterized 2D vector geometries and traditional deep CNN architectures [18, 19].
The geospatial community has adopted these approaches for a variety of remote sensing tasks, including vector shape generation. The Microsoft open building footprints dataset [20] was generated by segmenting and subsequently vectorizing building polygons based on satellite and aerial imagery using deep semantic segmentation networks [21].
Such expertise in image classification naturally influenced approaches to learning tasks on vector shapes. Xu et al. [22] trained a deep convolution autoencoder to assess the quality of geographic information of rasterized building footprints collected from OpenStreetMap (OSM) [23]. Veer et al. [2] evaluated the classification performance of deep neural networks on spatial vector geometries with geometry vertices encoded as 1D sequences directly feed into deep learning models. While this shows the ability of CNNs and RNNs to achieve performances comparable to machine learning methods with hand-crafted features, these approaches neglect the desirable property of learning transformation invariant features of shapes.
Beyond the geospatial community, the typography and computer graphics communities also significantly handle vector shapes. Lopes et al. [24] developed a generative model with convolution layers as encoders, to generate vector text glyphs. Yet, this model is optimized to learn only scale-invariant features of font styles. Mino and Spanakis [25] and Sage et al. [26] generated vector icons and logos from rasterized images using deep generative autoencoders proposed by Radford et al. [27] and Odena et al. [28]. Furthermore, Ha and Eck [29] proposed a recurrent neural network (RNN), sketch-rnn, to generate stroke-based vector drawings of common objects, while later Carlier et al. [30] introduced a hierarchical generative network (DeepSVG) which uses Transformer blocks [31] as the encoder and decoder of a backbone architecture for scalable vector graphics (SVG) generation.
2.2 Vector Geometry as Discrete Data Representation
Recently, the progress of graph representation learning [32, 33, 16, 34], enabled studies re-formalized learning tasks of vector polygons as graph or set representation learning problems. Previous work by Yan et al. [1] investigated the building pattern classification problem utilizing graph convolution networks. They processed buildings into building clusters, based on geometric patterns of individual buildings. They discovered the relational representations between buildings in graphs built using clusters based on Delaunay triangulation and Minimun Spanning Tree. They then trained a graph convolution model [16] for a binary classification of regular and irregular building clusters. Similarly, Bei et al. [15] devised a spatial adaptive model that makes use of graph encoding, but applied to entire building clusters (i.e., buildings as graph nodes) and applied a graph convolution neural network [16] for group pattern recognition (i.e., inner-state, edge-state and free state building patterns).
Most recently, Yan et al. [6] presented a graph convolutional autoencoder, based on the graph convolution model [16], to analyze the shape of building polygons. Different from previous works, they proposed to represent building polygons as graphs with the vertices of boundary geometries of buildings as graph nodes and their neighbour vertices as description features. However, this method did not consider the relational structures between connected vertices in graphs. Hence, their experiments on building shape coding and empirical results of building shape similarity measurement suggest that the proposed method is sensitive to rotation of building polygons.
Liu et al. [35] proposed a deep point convolutional network (DPCN), a modification of a DGCNN [36], to recognize the geometric shapes of buildings in map space, with the convolution operator TriangleConv. Their method operates on the coordinates of vertices of the exterior boundaries of building polygons and the convolution operator TriangleConv extracts local geometric features from triangle features, where triangle features are given by $f_{\verb+TriangleConv+}(x_{i-1},x_{i},x_{i+1})=\{(x_{i}-x_{i-1}),(x_{i}-x_{i+1}),(x_{i-1}-x_{i+1})\}$, and $(x_{i-1},x_{i},x_{i+1})$ are adjacent coordinate features of a building polygon. This method, trained and evaluated on a building footprint dataset [6], achieves fair classification performances (i.e., accuracy) on geometric shapes of buildings. However, their method requires a pre-definition of local triangles constructed by the graph nodes and adjacency points, and attempts to aggregate the local triangle features (i.e., angles and areas) for updating node features. We argue that learning additional local triangle features of polygons repeatedly is redundant and increase computing overhead in model training, as pointed out by Tailor et al. [37]. Further, they do not evaluate the robustness and generalizability of permutation-invariant data representations and learning architectures to geometric transformations of spatial geometries.
Mai et al. [8] recently proposed a general-purpose polygon encoding model for single and multi-part polygonal geometries. Their polygon encoding method, ResNet1D is a 1D CNN-based polygon encoder that captures local geometric features of spatial polygons and is deemed suitable for shape classification. ResNet1D encodes a polygon coordinate vertex $i$ as relative offset vectors from $k$ neighbouring vertices: $\{x_{i},x_{i-1}-x_{i},x_{i+1}-x_{i},\cdots,x_{i+k}-x_{i},x_{i-k}-x_{i}\}$. They contrast this encoding with a Non-Uniform Fourier Transformation (NUFT) method NUFTspec, based on the shape spectral domain. The NUFTspec aggregates global features of polygons as a NUFT representation and is a better encoding method for geometry topological relation predictions. Their work is thus parallel to our study. Here, we focus on the ability to support specific transformation invariant properties of combinations of learning architectures (permutation-variant CNN, set-based, graph-based and attention-based permutation-invariant models) and data representations (1D grid, set, graph and complete graph).
While some of the above studies leverage graph representations and local geometric encoding of vector polygons and state-of-art graph representation learning models to solve tasks on vector shapes, they do not yet provide a grounded, principled justification for graph-based learning methods on vector polygons. We attempt to fill this gap here by providing experimental results grounded in a sound theoretical motivation, and demonstrate why discrete data representations in combination with permutation-invariant learning models provide desired robustness to vector shape learning tasks.
2.3 Graph Representation of Polygon
Polygons are point sets connected by lines, constituting a collection of clock-wise or counter clock-wise linear rings [38]. According to the Gestalt principles of invariance, the compact embedding of linear rings should ideally be invariant to geometric transformations (rotation, scale and shear). We define a graph $\mathcal{G}=(X,E)$, where $X$ is a node matrix that contains the point coordinates of geometry vertices and $E$ is an edge matrix that captures the connectivity of these vertices along the linear rings (boundaries of the outside and of any holes of polygons), hence encoding the node connectivity of the graph. The binary adjacency value $a_{ij}=1\in A$ is used to indicate the connectivity between nodes $i$ and $j$, with $A$ being a square adjacency matrix.
2.4 Graph Representation Learning
Graph representation learning model leverages the property of data permutations and learns a whole graph embedding $\mathcal{G}^{\star}$ from the input graph $\mathcal{G}$ with differentiable permutation-invariant function $\mathcal{F}$. The learned embedding $\mathcal{G}^{\star}$ can be applied to graph-wise and node-wise classifications, or even edge predictions.
By representing polygon linear rings as graphs, we can define a differentiable permutation-invariant function $\mathcal{F}$, which takes the graph representations as inputs and returns a corresponding compact graph embedding $\mathcal{F}(\mathcal{G};\theta):=\mathcal{F}(X,A;\theta)\rightarrow\mathcal{G^{\star}}\in\mathcal{R}$.
Considering a permutation matrix $\mathcal{P}$ operating on graph $\mathcal{G}$, the permutation matrix $\mathcal{P}$ alters the order of node feature matrix $X$ in $\mathcal{G}$ along with the adjacency matrix $A$, producing a perturbed graph $\mathcal{G}^{\prime}=(PX,PAP^{\intercal})=(X^{\prime},A^{\prime})$. The ideal permutation-invariant function $\mathcal{F}$ should satisfy $\mathcal{F}(X^{\prime},A^{\prime};\theta)=\mathcal{F}(X,A;\theta)$. We relate the idea of permutation matrix $\mathcal{P}$ to the aforementioned geometric transformations.
In the view of invariant feature learning, the permutation invariance of $\mathcal{F}$ outputs are guaranteed through the local aggregations of node features of $x\in X$ and the linear transformations of aggregated node features:
$\mathcal{F}(X;\theta)=\rho(\sum_{x\in X}\phi(x,\theta)).$
The local differentiable function $\phi$ linearly transforms each node feature individually $x$ to latent space and the aggregation function sums over latent node features  [^1] The basic aggregation in this setting can also be averaging or max pooling., and $\rho$ is a global differentiable function applied on the summed node features followed by a non-linear activation function. In this simple setting, the outputs of $\mathcal{F}$ are invariant to the permutation of node features since the aggregation sum returns the same outputs for any input permutation: $Aggr_{sum}(x_{1},x_{2},....,x_{n})=Aggr_{sum}(x_{n},...,x_{2},x_{1})$.
This learning setting [9] makes a simple assumption on the permutations acting on input data, which can affect only individual node features without taking into account the connectivity within data (i.e., boundaries of polygon geometries). Graph convolution neural networks (GCN) by Kipf and Welling [16] learn convolutional features of node matrix $X$ with the normalized graph Laplacian matrix $\hat{A}=\tilde{D}^{-1/2}\tilde{A}\tilde{D}^{-1/2}$, where $\tilde{A}=A+I$ is the adjacency matrix with a identity matrix (i.e., self-connections of nodes), and $\tilde{D}$ is the diagonal degree matrix. Intuitively, the graph Laplacian matrix describes the divergence of energy flowing from one place (source nodes) to another (target nodes) in a graph. Linking this back to the linear rings of polygons and graph representation, the graph Laplacian conceptually shows similar energy divergence for nodes on the same rings(i.e., polygons with holes) and for nodes that are close to each other (i.e, polygons without holes).
Thus, GCN can learn graph Laplacian embedding of node feature with a differentiable neural network layer as follows:
$\begin{array}[]{c}\mathcal{F}_{conv}(X,A;\theta)=\tilde{D}^{-1/2}\tilde{A}\tilde{D}^{-1/2}X\theta\\ =\hat{A}X\theta.\end{array}$ (1)
We can express Eq.1 from a vector-wise view:
$f(x_{i};a_{ij},\theta)=\sum_{\begin{subarray}{c}a_{ij}\in A,\\ x\in X\end{subarray}}\phi(x_{i},x^{\prime}_{j},\theta),\text{where}\;x^{\prime}_{j}=\frac{\tilde{a}_{ij}}{\sqrt{\tilde{d}_{i}\times\tilde{d}_{j}}}x_{j},$ (2)
$\frac{\tilde{a}_{ij}}{\sqrt{\tilde{d}_{i}\times\tilde{d}_{j}}}$ corresponds to the edge weight between nodes $x_{i}$ and $x_{j}$ in the normalized graph Laplacian matrix $\hat{A}$. From Eq.2, the latent feature of target node $x_{i}$ depends on the relation with its neighbour node $x_{j}$ and constant edge weight $\frac{\tilde{a}_{ij}}{\sqrt{\tilde{d}_{i}\times\tilde{d}_{j}}}$. However, the constant edge weight between nodes $x_{i}$ and $x_{j}$ largely limits the expressivity of latent node features of polygonal graph since every node has a constant node degree ($d=2$).
As discussed in Yan et al. [6], graph autoencoders GCAE based on GCN are sensitive to orientation and rotation of polygons, and therefore current graph-based learning methods are not optimized for geospatial settings with polygonal geometries of arbitrary orientation, or conversely, arbitrary observer orientation. Here we propose to leverage the message-passing mechanism [17, 39] with the permutation-invariant function $\mathcal{F}$ to learn expressive and robust latent features of polygons.
2.5 Message-passing Neural Network
The principle of the message-passing mechanism is the aggregation of neighbour node features in graphs based on the “message” computed between the source and target nodes. We define the message-passing function operating on a single node as:
$\begin{array}[]{c}h_{i}=\phi(x_{i},msg(x_{i},x_{j}),\theta),\end{array}$ (3)
where $\{x_{i},x_{j}\}\in A=1$ and $msg(\cdot)$ is a message function with the source and target nodes as inputs. From Eq.3, the node feature of $x_{i}$ is updated with the computed message $(x_{i},x_{j})$, and together linearly transformed into latent space, producing the latent feature $h_{i}\in\mathcal{H}$.
Eq. 4 present the message-passing neural network in vector view as follow:
$\begin{array}[]{c}\mathcal{F}_{msg}(X,A;\theta)=\rho_{\theta}(readout(h_{i}|h_{i}\in\mathcal{H})).\end{array}$ (4)
A global readout function (i.e., mean or max pooling) summarizes the latent node-wise features ($\mathcal{H}$) and computes a global graph embedding for the input graph by non-linearly transforming the pooled feature.
Hence, we propose a Messpage-passing neural network for Polygon geometries, PolyMP, as shown in Fig. 1. The PolyMP considers the point coordinates of polygons as input node features and the exterior connectivity of polygons as input edges. The message-passing layers of PolyMP then compute the “message” for each node according to the graph adjacency. Following existing works on point clouds [36, 11] and polygons Liu et al. [35], we compute the message of relative positions of two connected nodes: $msg(x_{i},x_{j})=|x_{j}-x_{i}|$, which enable aggregating the local geometric features of polygons for updating node features. After message-passing layers, a global pooling layer is applied over the polygons with aggregated node features to generate a whole graph embedding, which is then used for the downstream shape classification of polygons.
Figure 1: Model architecture of PolyMP. PolyMP consists of two message-passing layer and a classification layer (Multi-layer perceptron).
3 Experiment
3.1 Datasets
3.1.1 Glyph Dataset
We introduce a synthetic dataset of highly variable geometric shapes for benchmarking the classification performance of learning models on vector polygons. The dataset consists of 26 Latin alphabet character glyph geometries (here also equal to semantic classes A to Z) of fonts gathered from an online source [40]. An equivalent approach to construct a rich dataset of geometric shapes has been used to test algorithms in computational geometry before [41].
We extract the boundaries of glyphs ( i.e., contour lines, as called in typography) for 1,413 sans serif and 1,002 serif fonts, to produce 2D simple polygon geometries compliant with OpenGeospatialConsortium [38]. Serif and sans serif fonts are the two main typographic font families. Serif glyphs have salient decorative strokes ending the main character strokes increasing legibility of body text, whereas sans serif glyphs have clean strokes suitable e.g., for the display of headers. Importantly, these minor variations do not alter the Gestalt of the shapes (as proven by reader’s ability to consistently read the symbols). Thus, the labels assigned to the symbols are stable and robust. We encode each polygon geometry into a fixed-size feature matrix $\in R^{n\times 3}$ where each feature vector $\in R^{3}$ includes the 2D coordinates of vertices $(x,y)$ as geometric feature and a binary feature $(0,1)$ to indicate the positions of 2D coordinates, whether on outer rings or holes of polygons. Examples of glyph geometry are displayed in Figure 2.
Figure 2: Glyph dataset samples. Left to right columns: A-shape to J-shape glyphs.
3.1.2 OSM Dataset
Latin alphabet glyph geometries share geometric similarity with building footprints. We use a building geometry dataset designed to test the generalizability of learning models on spatial objects proposed by Yan et al. [6]. The dataset contains 10,000 real-world building footprints extracted from OSM [23], labelled by 10 categories based on a template matching to letters [42]. The building footprints are randomly rotated and reflected from the original canonical samples (Figure 3). This is currently the most comprehensive real-world dataset of vector geometries with purely shape-based labels (as opposed to, e.g., building use, which should not be used for shape classification). We note that the assignment of labels in this dataset is somewhat subjective and likely less robust than in the glyph-based dataset. Note, in particular, that shape reflection is problematic for label stability, e.g., for asymmetric letters ($R$), or mirror shapes of distinct labels ($S$,$Z$).
Figure 3: OSM dataset samples. Top row: building geometries of standard shapes. Bottom row: geometric transformation counterparts. Left to right columns: E, F, H, I, L, O, T, U, Y and Z-shape buildings. [6]
3.1.3 Pre-processing
Label-preserving transformations
To examine the invariances of geometric transformations of polygon representations and corresponding learning models with respects to data transformations, we consider four training datasets under data transformations, each consisting of 20%, 40%, 60% and 80% of glyph polygons that have undergone label-preserving geometric transformations. The details of transformations are presented in Table 1. The original and transformed data samples are visualised in Fig. 4. We thus obtain five training datasets, including the original dataset: dataset 0%, 20%, 40%, 60% and 80% for the experiment. The test set consists of data randomly sampled from the five datasets.
Table 1: Label-preserving transformations applied to samples of Glyph dataset maintains the semantic information (label) of original polygons (i.e., 180 degrees rotation of letter M alter its semantic label, converting to letter W).
Transformation Operation
Rotation Rotate data around its centroid by a random angle $\in$ $\left[-75^{\circ},75^{\circ}\right]$.
Scaling Scale data by random distinct factors $\in$ $\left[0.1,2\right]$ on the $x$ and $y$ axes.
Shearing Shear data in a random angle $\in$ $\left[-45^{\circ},45^{\circ}\right]$ on the $x$ and $y$ axes.
Figure 4: Data samples of Glyph dataset under label-preserving transformations.
Polygon Simplifications
To test the invariances to trivial vertices of polygon, we apply the Douglas–Peucker algorithm [43] (tolerance = 1.0) on the original size of glyph polygon (in the order of 50 $\times$ 50 units) to generate simplified samples for glyph dataset. The resulting polygon linestrings are defined by fewer vertices (in particular, excluding all co-linear or close to co-linear vertices). Notice that the simplification occurs before the normalisation to the range of (-1,1) and any augmentation of polygons, ensuring that the simplified geometries preserve the topology and essence of the shape of the original glyph (see Fig. 5).
Figure 5: Visualisation of original and simplified polygon examples generated with Douglas–Peucker algorithm [43]. The simplified polygons, compared to original polygons, contain a subset of the vertices that defined the original curve.
3.2 Training Configuration
We evaluate the performances of learning models from previous studies, in particular: the VeerCNN in Veer et al. [2], based on 1D sequences converted from polygons as inputs; the DeepSet from Zaheer et al. [9] and SetTransformer Lee et al. [12], based on polygon vertices as input point clouds; and the GCAE in Yan et al. [6], representing polygons as input graphs. We then compare their performances with the proposed message-passing model for polygons (PolyMP).
The model configurations used in our experiments are in Table 2. Correspondingly, we depict the data representations of polygons and baseline models in Figure 6. The proposed PolyMP is demonstrated in Figure 1. For experimental control, every model tested was constrained to a comparable model depth and latent space of feature embedding.
For model training on Glyph dataset and fine-tuning on OSM dataset, we compute the cross entropy loss [44] to determine training gradients for fast training convergence and good model generalizability. We use the Adam optimization [45] with an initial learning rate of $0.01$. For better model generalization, we adopt a learning rate decay strategy that reduces the learning rate by a factor of $10$ once the model training reaches a performance plateau over $25$ epochs. An early-stopping mechanism halts the training once the loss ceases to decrease for $50$ epochs, to mitigate overfitting. We set the batch size to $64$ and the total training epochs to $100$. Thus, importantly, the models here are not designed to individually outperform current state-of-the-art for each architecture, but to maintain comparability across architectures and to enable evaluation of the effects of data representations on the results.
Table 2: Model configuration. Each model is configured to have the same number of convolution layers and dimensions of latent embedding, to keep the model size (i.e., number of training parameters) comparable.
Model Conv.
depth
Latent
dim.
#Param. Pooling
VeerCNN [2] 2 {2, 32, 64} 13,853 Global avg.
DeepSet [9] 2 {2, 64, 64} 11,837 Global avg.
SetTransformer [12] 2 {2, 64, 64} 113,531 Attention
GCAE [6] 2 {2, 64, 64} 7,613 Global avg.
PolyMP (Ours) 2 {2, 64, 64} 11,837 Global avg.
Figure 6: Data representations of polygons (as 1D sequence, Point Set, Graph and Complete Graph) and the corresponding learning models (Baselines: VeerCNN, DeepSet, GCAE and SetTransformer).
4 Result
We report model performances for all experiments as the overall accuracy (O.A.), calculated as the ratio of true positives over the total number of samples.
4.1 Results on Glyph Dataset
Table 3 and 4 demonstrates the model performances testing on Glyph dataset and the simplified polygons, respectively.
At $0\%$ transformation ratio in Table 3, we observe that the PolyMP achieves $99.41\%$, $42.91\%$ and $80.03\%$ acc. on the original, rotated and scaled test samples, whereas VeerCNN achieves $59.07\%$ acc. on sheared samples. While deepset, set transformer and GCAE maintain comparable performances ($94.62\%$, $65.01\%$; and $96.25\%$, $67.91\%$; and $95.16\%$, $64.54\%$, respectively) on original and scaled test samples, the models suffer large performance deterioration, achieving only $23.67\%$, $26.00\%$ and $25.23\%$, $27.79\%$ and $31.38\%$ and $38.84\%$ test acc. on rotated and sheared samples, respectively. From the overall accuracy, the VeerCNN outperforms the PolyPM by a slight margin ($69.43\%$ and $69.01\%$).
At 20% transformation ratio, the PolyMP improves model test acc. on rotated, scaled and sheared samples rapidly, achieving 83.17% (+40.26%), 91.29% (+11.26%) and 86.86% (+33.54%) respectively. In overall acc., the PolyMP achieves 90.12% O.A. (+21.11%) and outperforms VeerCNN. In comparison, the other models can not match with PolyMP and achieve comparable performances from increasing data diversity of training set.
At 40% and 60% transformation ratio, we notice that the accuracy improvements gained from increasing the ratio of geometric transformed samples in training set have converged. In this case, the PolyMP maintains its leading performances among models (91.91% and 93.06% O.A.), and VeerCNN (83.30% and 84.88% O.A.), deepset (80.12% and 81.49% O.A.), set transformer (80.33% and 81.94% O.A.) and GCAE (72.72% and 75.50% O.A.) fall behind.
At 80% transformation ratio, the PolyMP achieves the optimal test performances at rotated and sheared samples, giving 90.35% and 91.52% test accuracy. This empirical results suggest that compared to the sequence encoding of polygons and CNN as well as the set representation and set-based learning models, the graph representation together with message-passing neural network (i.e., PolyMP) are robust to the permutations of feed-in order of polygon vertices, which occurs in geometric transmutations of polygons, in particular rotation and shearing.
From a different perspective, the ratio of original samples is decreasing in the training set progressively from the 0% to 80% transformation ratios datasets. The PolyMP only encounters 1.46% test performance deterioration along this progression (99.41% to 97.95% acc.), while the VeerCNN (-4.13% acc.), deepset (-3.96% acc.), set transformer (-5.08% acc.) and GCAE (-5.58% acc.) suffer larger performance drops on the test accuracy of original samples.
Table 3: Test performances on Glyph dataset. Transformation ratios 0%, 20%, 40%, 60% and 80% refer to training datasets consisting of 0%, 20%, 40%, 60% and 80% geometric transformed samples. Glyph-O, Glyph-R, Glyph-SC and Glyph-SH Acc. indicate the classification accuracy of models tested on original, rotated, scaled and sheared data. Overall accuracy is reported as Glyph O.A..
Trans. Ratio Model Glyph-O Acc. Glyph-R Acc. Glyph-SC Acc. Glyph-SH Acc. Glyph O.A.
0% VeerCNN 97.64 41.51 79.15 59.07 69.43
DeepSet 94.62 23.67 65.01 26.00 52.56
SetTransformer 96.25 25.23 67.91 27.79 54.42
GCAE 95.16 31.38 64.54 38.84 57.83
PolyMP 99.41 42.91 80.03 53.32 69.01
20% VeerCNN 95.12 67.63 83.53 77.78 81.05
DeepSet 93.88 67.95 77.31 64.72 76.25
SetTransformer 93.87 71.72 82.43 72.56 80.18
GCAE 93.31 47.31 71.47 57.11 67.61
PolyMP 98.92 83.17 91.29 86.86 90.12
40% VeerCNN 94.75 73.26 84.77 80.29 83.30
DeepSet 92.57 74.24 79.24 72.08 80.12
SetTransformer 91.97 72.73 83.25 73.25 80.33
GCAE 91.70 59.55 74.35 64.42 72.72
PolyMP 98.48 87.06 92.49 89.25 91.91
60% VeerCNN 94.50 77.16 85.38 82.36 84.88
DeepSet 91.72 76.89 81.46 73.97 81.49
SetTransformer 90.93 77.04 82.39 77.31 81.94
GCAE 90.46 65.02 76.38 68.78 75.50
PolyMP 98.30 89.40 93.43 91.11 93.06
80% VeerCNN 93.51 78.97 85.25 83.17 85.25
DeepSet 90.66 78.68 82.14 76.55 82.47
SetTransformer 91.17 79.94 83.69 79.78 83.67
GCAE 89.58 67.95 77.99 70.55 76.83
PolyMP 97.95 90.35 93.30 91.52 93.42
In Table 4, the test performances of models trained and evaluated on the simplified polygons of Glyph dataset are presented. In this experiment setting, we focus on models’ test performances on simplified polygons containing fewer numbers of vertices, and compare them with the test performances showed in Table 3.
We observe similar test performance trends to those of models evaluated on the simplified polygons. Noticeably, there is a significant performance deterioration in VeerCNN compared with the test performances in Table 3 on simplified polygons. Specifically, at 0% transformation ratio, the test performances of VeerCNN drops from 69.43% O.A. (Table 3) to 66.56% O.A. (Table 4), i.e., a 2.87% performance deterioration. As the data diversity increases, the performance deterioration in VeerCNN increases, too, from 5.2%, progressively through to 5.56% O.A. deterioration respectively from 20% to 80% transformation ratios.
The performance deterioration of VeerCNN strongly indicates that CNN models are not robust to varying the numbers of vertices in polygons. This can be attributed to a number of factors, incl. the data encoding of polygons in VeerCNN as as fixed-length 1D sequences with zero padding. Simplifying polygons and removing trivial vertices from polygons alters the length of 1D sequences, and hence the geometric information of simplified polygons are not only diluted by the zero padding in the 1D sequences, but also by alteration of vertex adjacency explicitly encoded in the 1D sequence.
In comparison, we notice only small performance drops occur in general set-based and graph-based learning models. We credit this to set and graph representation of polygons and the corresponding learning models as both set and graph-based models accept input polygons of varying lengths. In particular, the graph-based message-passing network (i.e., PolyMP) learns salient local node features from connected neighbours in the message-passing function with permutation-invariant pooling operations (i.e., Max-pooling), as shown in E.q. 4. A similar pooling operation is also adopted in set-based learning models [9, 12].
Table 4: Test performances on Glyph dataset, consisting of simplified polygons generated by Douglas–Peucker algorithm (tolerant=1.0) with the removal of trivial vertices.
Trans. Ratio Model Glyph-O Acc. Glyph-R Acc. Glyph-SC Acc. Glyph-SH Acc. Glyph O.A.
0% VeerCNN 94.72 39.65 75.42 56.12 66.56
DeepSet 94.01 23.75 65.11 25.53 52.30
SetTransformer 94.57 23.67 65.72 26.61 52.76
GCAE 94.83 31.64 64.21 39.90 58.08
PolyMP 98.89 43.71 79.15 52.70 68.63
20% VeerCNN 90.55 62.58 77.80 72.31 75.85
DeepSet 93.81 68.26 77.83 64.93 76.47
SetTransformer 92.31 70.51 80.54 70.65 78.54
GCAE 92.26 47.29 71.28 57.33 67.39
PolyMP 97.61 82.45 88.42 84.84 88.22
40% VeerCNN 89.63 67.82 78.72 74.61 77.73
DeepSet 92.66 74.81 79.60 71.62 80.46
SetTransformer 89.86 70.91 80.70 70.97 78.14
GCAE 91.65 59.30 74.95 64.79 72.94
PolyMP 97.55 86.14 89.99 87.29 90.21
60% VeerCNN 89.29 71.91 79.42 77.29 79.50
DeepSet 91.58 77.44 81.62 73.84 81.57
SetTransformer 89.70 75.03 80.26 75.53 80.15
GCAE 90.06 63.61 76.64 68.05 74.79
PolyMP 97.35 88.73 91.04 89.31 91.47
80% VeerCNN 88.05 73.89 79.55 77.17 79.69
DeepSet 91.04 79.21 82.44 76.45 82.67
SetTransformer 89.99 79.14 82.70 78.60 82.63
GCAE 88.22 66.55 78.01 70.44 76.06
PolyMP 97.01 90.19 91.53 90.31 92.22
4.2 Results on OSM Dataset
We fine-tune models pre-trained on the glyph 80% dataset with the OSM dataset, and report the classification test accuracy of fine-tuned models in Table 5 for building footprints with original and normalised coordinates respectively.
PolyMP achieves optimal test performance on normalised footprints with $91.92\%$ O.A. on the OSM dataset, and performs better than all other methods. Both set-based methods (i.e., DeepSet: $95.71\%$ and SetTransformer: $94.99\%$) and graph-based GCAE ($94.79\%$) achieve comparable performance on OSM-O dataset with building footprints without geometric transformations. Conversely, the convolutional VeerCNN shows lower performance of $90.54\%$. When testing on OSM-R with randomly rotated and reflected footprints, PolyMP maintains a robust performance of $85.62\%$ and the baseline methods suffer performance deterioration at various levels (VeerCNN: $90.54\%\rightarrow 78.51\%$, DeepSet: $95.71\%\rightarrow 84.86\%$, SetTransformer: $94.99\%\rightarrow 83.87\%$, GCAE: $94.79\%\rightarrow 80.19\%$).
Performances of models fine-tuned on OSM footprints with original spatial coordinates drop for all models except for PolyMP, which maintains a performance of $81.58\%$ O.A.. This robust performance of PolyMP against the translation of spatial coordinates can be accounted for by the message-passing network of PolyMP where the local geometric features of polygons are computed (i.e., the relative positions of polygons $msg(x_{i},x_{j})=|x_{j}-x_{i}|$). This relative positional features reduce the impact of absolute positioning and effectively normalise the spatial coordinates of polygons while preserving the geometric properties of polygons. VeerCNN outperforms DeepSet ($24.34\%$ O.A) and CGAE ($23.18\%$ O.A) and achieves a test performance of $42.68\%$ O.A. when fine-tuned on the original OSM footprints. We account for VeerCNN’s performance by the built-in translation invariant priors in conventional convolution neural networks. Similarly, SetTransformer shows comparable performance as VeerCNN ($44.72\%$ O.A), which can be attributed to the self-attention features based on the fully connected graph of point sets.
In Fig. 7, we visualise the classification results of PolyMP evaluated on spatial building footprints in OSM-O (Fig. 7 (a)) and OSM-R (Fig. 7 (b)). Predictions indicate that the fine-tuned PolyMP can robustly classify building footprints on a map in varying sizes and geometric shapes based on the geometric invariant features learnt by message-passing networks, taking the graph-based representation of polygons as inputs.
Table 5: Fine-tune test performances on OSM dataset. OSM-O and OSM-R acc. suggest the the classification accuracy of models tested on original and rotated & reflected samples. Overall accuracy is reported for OSM O.A..
Model OSM-O Acc. OSM-R Acc. OSM O.A.
VeerCNN 90.54 78.51 84.52 Normalised coordinates $\in(-1,1)$
DeepSet 95.71 84.86 90.56
SetTransformer 94.99 83.87 89.42
GCAE 94.79 80.19 88.02
PolyMP 98.04 85.62 91.92
VeerCNN 42.35 43.01 42.68 Original coordinates
DeepSet 24.12 24.08 24.34
SetTransformer 44.67 44.77 44.72
GCAE 23.44 23.04 23.18
PolyMP 89.86 73.56 81.58
(a) OSM-O
(b) OSM-R
Figure 7: Shape classification of building footprints by PolyMP on the selected region in OSM dataset (downtown areas of Shanghai, China [6]). Model predictions are colored in red.
5 Discussion
(a) Glyph dataset.
(b) OSM dataset.
Figure 8: Failure cases of PolyMP on Glyph and OSM datasets.
(a) Sheared M glyph.
(b) Rotated M glyph.
(c) Scaled W glyph.
(d) Rotated W glyph.
(e) Sheared R glyph.
(f) Original Q glyph.
Figure 9: Feature maps to visualize salient features learned by individual models to classify input geometries to output categories. The darker the color, the more salient are the features learned. Feature values are normalised $\in[0,1]$.
5.1 Model Expressivity
For the interpretation of model expressivity with respect to various polygon encodings and learning models, we visualize the feature maps of models tested on original and geometric transformed samples in glyph dataset in Figure 9.
Figs. 9a and 9b depict the salient feature maps and class predictions of four models (Deepset, Set Transformer, GCAE and PolyMP) on geometric transformed samples (sheared and rotated glyph M). In this case, the PolyMP correctly predicts the class label of glyph $M$, whereas DeepSet, Set Transformer and GCAE have made incorrect predictions on the geometric transformed samples. By comparing the feature maps in Figures 9c and 9d, the PolyMP learns to identify the global structures or the “skeleton” of polygons via non-trivial vertices on the boundary. The message-passing network and local pooling function (i.e., max-pooling) can capture the salient vertices from local neighbours (i.e., non-trivial vertices (dark color) in the clusters of trivial vertices (light color), which enable a robust embedding of polygons under geometric transformations. Set-based models and GCAE learn sparser geometric features for input polygons, capturing salient features regionally in node clusters and ignoring geometric features of more distant nodes. This can be attributed to the set representation where any relational information between nodes is neglected.
Figures 9e and 9f demonstrate the feature maps on polygons with holes. Both Deepset and Set Transformer only capture geometric features from non-trivial vertices located at the exterior boundary of polygons. While set-based models accept point set encoding of polygons in varying length and the output features are invariant to the feed-in order of vertices, such methods largely ignore the connectives and interactions between points. Such a learning setting inherently hinders the model’s capability to learn geometric information from polygons with holes. The graph convolution model, GCAE, lacks learning flexibility in learning local geometric features as discussed in the previous sections. On the other hand, PolyMP can capture geometric features from salient points located at the exterior and interiors of polygons. In the message-passing network, the exterior and interior of polygons are considered to be two distinct elements as there is no direct connection between points located at the two boundaries. In other words, the geometric features (i.e., the “message”) are only passed between points that are connected in a connected component of a graph, in this case defined by the linear rings of polygon boundaries.
5.2 Limitations
Of course, PolyMP still encounters limitations in terms of misclassification. Fig. 8. demonstrates failures of PolyMP classifications on the Glyph and OSM datasets. On synthetic data, misclassifications occur predominantly on samples where excessive geometric transformations (e.g., heavy shearing and anisotropic scaling) destroy the gestalt of the shape to the extent it is no longer label preserving (Fig. 8(a): letter V, Z and R). A special case are shapes with different labels dependent on orientation – M and W, where depending on the extent of rotation and shearing, misclassifications occur. For instance in row two of Fig. 8(a), we observe that rotated letters W can be misclassified into letter M, or even Z; heavily sheared W can be misclassified as H and V. Similarly for building polygons (OSM), PolyMP may misclassify polygons sharing similar geometric shapes (E, vs L or U) (Fig. 8).
Considering the characteristics of the two datasets (Sec. 3.1), PolyMP trained Glyph dataset learns locally salient geometric features via message-passing layers from serif glyphs, which contain salient decorative strokes ending. Letters sharing similar geometric shapes are misclassified when the geometric shapes of classified polygons are destroyed by excessive geometric transformations.
6 Conclusion
Based on extensive empirical evidence given by the analysis of invariances to geometric transformations, trivial vertices and model generalizability, we demonstrate that discrete, non-grid like data representations in combination with graph-based message-passing neural networks (PolyMP) are optimal for deep learning models on 2D vector geometries. We demonstrate a robust ability to learn shapes of polygonal geometries, and demonstrate the ability to generalize to a real-world building footprint dataset.
We find that a learning model based on graph data representation that explicitly encodes both node and explicit edge structures leads to a robust approach enabling highly accurate feature learning on geometric shapes – conveniently, these structures are readily available from the 2D simple feature representations of geometries implemented in current GIS libraries. Our results indicate that representations play an important part providing a strong inference bias for geometric shape learning, and tasks such as the classification of polygons and boosting model performance. We believe that our approach to learning geometric-invariant features of polygons is not limited to simple geometric shape classification. This approach can benefit suitable geospatial tasks where the geometric-invariant features of polygons matches the learning objectives (i.e., geospatial regression tasks involving spatial relation predictions of polygon map [8]).
We also discuss the limitations of PolyMP through misclassification examples on datasets. Through the failure cases of classification, we observe that the message-passing network for graph data representation of polygons has limited capability of transferring locally salient geometric features from one domain to another. The simple message-passing network has limited control on learning/not learning geometric invariant features as need.
The geometric invariant feature learning of vector geometries is currently an area of intense investigation in spatial deep learning [2, 1, 6, 35, 8]. Unlike in learning based on raster datasets or point clouds, the segmentation problem is partially solved by the explicit encoding of structured vector geometries. Here, we show results that allow us to undertake the classification of spatial geometric shapes efficiently, based on decisions that can be grounded in theory and are experimentally validated. In future work, we will further examine how richer graph encoding of the structure of geometries, as well as of the relationships between proximal geometries could further boost model performance, or enable the segmentation of complex spatial object arrangements.
Statements and Declarations
Data availability. The data and codes that support the findings of this study are available with the identifier(s) at the private link: https://figshare.com/s/b781d7724d3220e24c7c.
Conflicts of interest. This paper has been approved by all co-authors. The authors have no competing interests to declare that are relevant to the content of this article.

### v2

@@
 Abstract
-Geometric shape classification of vector polygons remains a non-trivial learning task in spatial analysis.
-Previous studies mainly focus on devising deep learning approaches for representation learning of rasterized vector polygons, whereas the study of discrete representations of polygons and subsequent deep learning approaches have not been fully investigated.
-In this study, we investigate a graph representation of vector polygons and propose a novel graph message-passing neural network (PolyMP) to learn the geometric-invariant features for shape classification of polygons.
-Through extensive experiments, we show that the graph representation of polygons combined with a permutation-invariant graph message-passing neural network achieves highly robust performances on benchmark datasets (i.e., synthetic glyph and real-world building footprint datasets) as compared to baseline methods.
-We demonstrate that the proposed graph-based PolyMP network enables the learning of expressive geometric features invariant to geometric transformations of polygons (i.e., translation, rotation, scaling and shearing) and is robust to trivial vertex removals of polygons.
-We further show the strong generalizability of PolyMP, which enables generalizing the learned geometric features from the synthetic glyph polygons to the real-world building footprints.
+Geometric shape classification of vector polygons remains a challenging task in spatial analysis.
+Previous studies have primarily focused on deep learning approaches for rasterized vector polygons, while the study of discrete polygon representations and corresponding learning methods remains underexplored.
+In this study, we investigate a graph-based representation of vector polygons and propose a simple graph message-passing framework, PolyMP, along with its densely self-connected variant, PolyMP-DSC, to learn more expressive and robust latent representations of polygons.
+This framework hierarchically captures self-looped graph information and learns geometric-invariant features for polygon shape classification.
+Through extensive experiments, we demonstrate that combining a permutation-invariant graph message-passing neural network with a densely self-connected mechanism achieves robust performance on benchmark datasets, including synthetic glyphs and real-world building footprints, outperforming several baseline methods.
+Our findings indicate that PolyMP and PolyMP-DSC effectively capture expressive geometric features that remain invariant under common transformations, such as translation, rotation, scaling, and shearing, while also being robust to trivial vertex removals.
+Furthermore, we highlight the strong generalization ability of the proposed approach, enabling the transfer of learned geometric features from synthetic glyph polygons to real-world building footprints.
 keywords:
@@
 The ability to reflect Gestalt principles through transformation invariances thus present a strong motivation for an inductive bias in the design of learning-based models for geometric shape recognition of spatial objects.
-Spatial objects are often conveniently represented as vector polygons, a data representation thus far neglected in deep learning research.
+Spatial objects are often conveniently represented as vector polygons, a discrete data representation thus far neglected in deep learning research.
 Learning geometric-invariant features from vector polygons has the following requirements: (1) a generic data representation that encodes geometric features of polygons without information loss; and (2) a learning model built on this input data representation that enables learning latent geometric-invariant features robust to geometric transformations.
@@
 [8] as “[vertices] …where the addition or removal of the vertex have no effect on the geometric shape and topological properties of the outlines of the polygons”.
-In this study, we treat the addition or removal of trivial vertices of polygons as changes that do not alter the semantic information of the shape (captured by the human-assigned label), thus closely reflecting the relationship with the Gestalt principles.
+The addition or removal of trivial vertices in polygons does not alter the semantic information of the shape (as captured by the human-assigned label), closely aligning with the Gestalt principles.
 Learning models tailored for discrete (i.e., not grid-like), permutation invariant data representations are predicated on data domains where entries hold no explicit neighborhood information (e.g., point sets [9, 10, 11, 12]).
@@
 Graph-based representations also enable the feed-in order of polygon vertices to be independent of the model outcomes, while the connectivity (topology) between parts of polygon representations remains invariant to geometric transformations.
-Graph convolutional autoencoders (GCAE) [6] extend graph convolution neural networks [16] by learning spectral convolution embeddings from polygon graphs and demonstrate the effectiveness of latent embedding.
+Graph convolutional autoencoders (GCAE) [6] extend graph convolutional neural networks (GCNs) [16] by learning spectral graph embeddings from polygon graphs, demonstrating the effectiveness of graph latent embeddings in polygon shape retrieval.
 Bronstein et al.
-[13] discussed that the graph convolution layers aggregate node features from neighboring nodes with constant weights, largely limited by the topological structure of the input graphs.
-In vector polygons vertices on the linear rings have constant neighbours (i.e., left and right neighbour vertices on linear rings), leading to a reduction in the expressivity of graph convolution features learned from polygon graphs.
-In this study, we propose to leverage graph message-passing mechanisms [17] in graph-based learning models to learn highly expressive and robust latent features of polygons.
-Our hypothesis is that graph representations of polygons enable the learning of robust latent features of polygons invariant to geometric transformations (i.e., rotation, scaling and shearing) when combined with graph message-passing models, thus supporting generalizability on data with domain shifts.
-If the hypothesis holds, we suggest that the downstream tasks performed on the learned features of vector polygons lead to robust performance on shapes with varying amounts of trivial vertices.
-We present the results of a series of experiments designed to evaluate model robustness on polygons altered by geometric transformations, and investigate the generalizability of the findings on combinations of polygon representations and model architectures (permutation-invariant vs.
-translation-invariant models) for the classification of vector polygons with and without holes (inner linear rings).
-After a literature review of related literature (Section 2) we present a synthetic dataset of polygonal shapes of high shape variability with known, robust human labels, based on latin alphabet character glyphs (Section 3.1.1).
-Despite a broad range of shape variations between fonts (in particular with respect to trivial vertices), characters must maintain strong recognisability attributable to the shapes’ Gestalt.
-We then experimentally benchmark the performance of learning models operating on three distinct data representations (1D sequence, set and graph) and varying amounts of geometric transformations (i.e., rotation, scaling and shearing).
-Our results (Section 4) document the models’ robustness to geometric transformations by benchmarking their performance on the synthetic glyph dataset.
-Building on these results, we further evaluate the generalizability of the models on a real-world building footprint dataset collected from the Open Street Map (OSM) [6].
+[13] noted that spectral graph convolutions aggregate node features from neighboring nodes with constant weights, making them highly dependent on the topological structure of the input graphs.
+In vector polygons, vertices on linear rings have fixed neighbors (i.e., the left and right adjacent vertices), which reduces the expressivity of graph convolutional features learned from polygon graphs.
+In this study, we leverage graph message-passing mechanisms [17] and propose a simple graph message-passing network, PolyMP, along with its densely self-connected variant, PolyMP-DSC, to learn more expressive and robust latent representations of polygons.
+We hypothesize that combining graph-based representations of polygons with graph message-passing models facilitates learning robust latent features that remain invariant to geometric transformations (i.e., rotation, scaling, and shearing), thereby improving the generalizability of shape classification across datasets.
+If our hypothesis holds, we suggest that downstream tasks performed on the learned features of vector polygons will exhibit robust performance on shapes with varying amounts of trivial vertices.
+We present a series of experiments designed to evaluate model robustness on polygons subjected to geometric transformations and investigate the generalizability of our findings across different combinations of polygon representations and model architectures (permutation-invariant vs.
+translation-invariant models) for classifying vector polygons with and without holes (inner linear rings).
+Following a literature review (Section 2), we introduce a synthetic dataset of polygonal shapes with high shape variability and well-defined human labels, derived from Latin alphabet character glyphs (Section 4.1).
+Despite significant shape variations across fonts, particularly in terms of trivial vertices, characters maintain strong recognizability due to their Gestalt properties.
+We then benchmark the performance of learning models on three distinct data representations (1D sequence, set, and graph) under varying geometric transformations (i.e., rotation, scaling, and shearing).
+Our results (Section 5) document the models’ robustness to geometric transformations through performance evaluation on the synthetic glyph dataset.
+Building on these findings, we further assess model generalizability using a real-world building footprint dataset from OpenStreetMap (OSM) [6].
 Our main contributions are:
   1. 1.
-    a thorough investigation of the impact of distinct combinations of graph representation of polygons and message-passing learning architectures for the learning of geometric features robust to geometric transformations of shapes and removals of trivial vertices, grounded in a theoretically grounded set of desiderata;
+    we provide a systematic investigation of vector polygon representations for geometric shape classification, highlighting the strengths of graph-based encoding over sequences and point sets;
   2. 2.
-    the contribution of an open synthetic dataset of vector polygon shapes based on glyphs, enabling the benchmarking of geometric learning models; and
+    the introduction of PolyMP and PolyMP-DSC, two lightweight graph message-passing architectures designed to learn geometric-invariant and robust latent representations;
   3. 3.
-    the experimental demonstration that graph representations of vector polygons in conjunction with message-passing neural networks achieve the most robust performance and generalization to real-world shape datasets.
+    the release of a synthetic dataset of vector polygons derived from character glyphs, enabling controlled benchmarking of geometric learning models; and
+  4. 4.
+    extensive experiments which show that the proposed message-passing models, PolyMP and PolyMP-DSC, significantly improve robustness to transformations and generalize well to real-world geospatial data.
+Together, these contributions aim to provide a principled and practical foundation for learning geometric-invariant representations of spatial polygons—pushing forward the capabilities of shape-based analysis in geospatial domains.
 2 Background
 2.1 Machine Learning with Vector Geometries
-In a straightforward extension of computer vision approaches devised for image classification, current research on polygonal shape classification typically learns salient geometric features of shapes based on rasterized 2D vector geometries and traditional deep CNN architectures [18, 19].
-The geospatial community has adopted these approaches for a variety of remote sensing tasks, including vector shape generation.
-The Microsoft open building footprints dataset [20] was generated by segmenting and subsequently vectorizing building polygons based on satellite and aerial imagery using deep semantic segmentation networks [21].
-Such expertise in image classification naturally influenced approaches to learning tasks on vector shapes.
+Building on computer vision approaches for image classification, polygonal shape classification has traditionally relied on learning geometric features from rasterized 2D vector geometries using deep CNN architectures [18, 19].
+The geospatial community has adopted these methods for various remote sensing tasks, including vector shape generation.
+For example, the Microsoft Open Building Footprints dataset [20] was generated by segmenting and vectorizing building polygons from satellite and aerial imagery using deep semantic segmentation networks [21].
+This expertise in image-based learning naturally influenced vector shape learning.
 Xu et al.
-[22] trained a deep convolution autoencoder to assess the quality of geographic information of rasterized building footprints collected from OpenStreetMap (OSM) [23].
-Veer et al.
-[2] evaluated the classification performance of deep neural networks on spatial vector geometries with geometry vertices encoded as 1D sequences directly feed into deep learning models.
-While this shows the ability of CNNs and RNNs to achieve performances comparable to machine learning methods with hand-crafted features, these approaches neglect the desirable property of learning transformation invariant features of shapes.
-Beyond the geospatial community, the typography and computer graphics communities also significantly handle vector shapes.
+[22] trained a deep convolutional autoencoder to assess the quality of rasterized building footprints collected from OpenStreetMap (OSM) [23].
+Similarly, Veer et al.
+[2] evaluated deep neural networks for classifying spatial vector geometries, encoding geometry vertices as 1D sequences directly fed into deep models.
+While CNNs and RNNs performed comparably to hand-crafted feature-based methods, they lacked the ability to learn transformation-invariant shape features.
+Beyond geospatial applications, typography and computer graphics research also leverage vector shape learning.
 Lopes et al.
-[24] developed a generative model with convolution layers as encoders, to generate vector text glyphs.
-Yet, this model is optimized to learn only scale-invariant features of font styles.
+[24] developed a convolution-based generative model to create vector text glyphs, but it was optimized primarily for scale-invariant font style learning.
 Mino and Spanakis [25] and Sage et al.
-[26] generated vector icons and logos from rasterized images using deep generative autoencoders proposed by Radford et al.
-[27] and Odena et al.
-[28].
-Furthermore, Ha and Eck [29] proposed a recurrent neural network (RNN), sketch-rnn, to generate stroke-based vector drawings of common objects, while later Carlier et al.
-[30] introduced a hierarchical generative network (DeepSVG) which uses Transformer blocks [31] as the encoder and decoder of a backbone architecture for scalable vector graphics (SVG) generation.
+[26] generated vector icons and logos from rasterized images using deep generative autoencoders [27, 28].
+Additionally, Ha and Eck [29] introduced Sketch-RNN, a recurrent neural network for generating stroke-based vector drawings, while Carlier et al.
+[30] proposed DeepSVG, a hierarchical generative model using Transformer-based encoding for scalable vector graphics (SVG).
 2.2 Vector Geometry as Discrete Data Representation
-Recently, the progress of graph representation learning [32, 33, 16, 34], enabled studies re-formalized learning tasks of vector polygons as graph or set representation learning problems.
-Previous work by Yan et al.
-[1] investigated the building pattern classification problem utilizing graph convolution networks.
-They processed buildings into building clusters, based on geometric patterns of individual buildings.
-They discovered the relational representations between buildings in graphs built using clusters based on Delaunay triangulation and Minimun Spanning Tree.
-They then trained a graph convolution model [16] for a binary classification of regular and irregular building clusters.
+Recent advances in graph representation learning [31, 32, 16, 33] have reformulated vector polygon learning as a graph-based or set-based problem.
+Yan et al.
+[1] applied graph convolution networks (GCNs) to building pattern classification, structuring buildings into clusters based on geometric relationships using Delaunay triangulation and Minimum Spanning Trees.
+Their GCN model classified building clusters into regular or irregular patterns.
 Similarly, Bei et al.
-[15] devised a spatial adaptive model that makes use of graph encoding, but applied to entire building clusters (i.e., buildings as graph nodes) and applied a graph convolution neural network [16] for group pattern recognition (i.e., inner-state, edge-state and free state building patterns).
-Most recently, Yan et al.
-[6] presented a graph convolutional autoencoder, based on the graph convolution model [16], to analyze the shape of building polygons.
-Different from previous works, they proposed to represent building polygons as graphs with the vertices of boundary geometries of buildings as graph nodes and their neighbour vertices as description features.
-However, this method did not consider the relational structures between connected vertices in graphs.
-Hence, their experiments on building shape coding and empirical results of building shape similarity measurement suggest that the proposed method is sensitive to rotation of building polygons.
+[15] introduced a spatially adaptive model using graph encoding for group pattern recognition, where buildings served as nodes in a graph convolutional network (GCN).
+Yan et al.
+[6] proposed a graph convolutional autoencoder (GCAE) for building shape analysis, encoding building polygons as graphs where boundary vertices acted as nodes.
+However, their method was sensitive to rotations, limiting its robustness for polygon shape retrieval.
+Addressing this limitation, Huang et al.
+[34] introduced a contrastive graph autoencoder to enable robust polygon shape matching and retrieval in large-scale vector datasets.
 Liu et al.
-[35] proposed a deep point convolutional network (DPCN), a modification of a DGCNN [36], to recognize the geometric shapes of buildings in map space, with the convolution operator TriangleConv.
-Their method operates on the coordinates of vertices of the exterior boundaries of building polygons and the convolution operator TriangleConv extracts local geometric features from triangle features, where triangle features are given by $f_{\verb+TriangleConv+}(x_{i-1},x_{i},x_{i+1})=\{(x_{i}-x_{i-1}),(x_{i}-x_{i+1}),(x_{i-1}-x_{i+1})\}$, and $(x_{i-1},x_{i},x_{i+1})$ are adjacent coordinate features of a building polygon.
-This method, trained and evaluated on a building footprint dataset [6], achieves fair classification performances (i.e., accuracy) on geometric shapes of buildings.
-However, their method requires a pre-definition of local triangles constructed by the graph nodes and adjacency points, and attempts to aggregate the local triangle features (i.e., angles and areas) for updating node features.
-We argue that learning additional local triangle features of polygons repeatedly is redundant and increase computing overhead in model training, as pointed out by Tailor et al.
-[37].
-Further, they do not evaluate the robustness and generalizability of permutation-invariant data representations and learning architectures to geometric transformations of spatial geometries.
+[35] developed a deep point convolutional network (DPCN), a modification of a DGCNN [36], for building shape recognition in map space.
+Their model introduced the TriangleConv operator, which extracts local geometric features from triangle-based representations of polygon vertices: $f_{\texttt{TriangleConv}}(x_{i-1},x_{i},x_{i+1})=\{(x_{i}-x_{i-1}),(x_{i}-x_{i+1}),(x_{i-1}-x_{i+1})\}$, where $(x_{i-1},x_{i},x_{i+1})$ are adjacent point features of a building polygon.
+Despite achieving competitive classification performance on building footprints [6], this method requires pre-defining local triangles of polygons and repeatedly computing redundant local geometric features (i.e., angles and areas) of polygons.
+Furthermore, it does not evaluate the robustness of learned representations to geometric transformations such as rotation, scaling, and vertex perturbations.
 Mai et al.
-[8] recently proposed a general-purpose polygon encoding model for single and multi-part polygonal geometries.
-Their polygon encoding method, ResNet1D is a 1D CNN-based polygon encoder that captures local geometric features of spatial polygons and is deemed suitable for shape classification.
-ResNet1D encodes a polygon coordinate vertex $i$ as relative offset vectors from $k$ neighbouring vertices: $\{x_{i},x_{i-1}-x_{i},x_{i+1}-x_{i},\cdots,x_{i+k}-x_{i},x_{i-k}-x_{i}\}$.
-They contrast this encoding with a Non-Uniform Fourier Transformation (NUFT) method NUFTspec, based on the shape spectral domain.
-The NUFTspec aggregates global features of polygons as a NUFT representation and is a better encoding method for geometry topological relation predictions.
-Their work is thus parallel to our study.
-Here, we focus on the ability to support specific transformation invariant properties of combinations of learning architectures (permutation-variant CNN, set-based, graph-based and attention-based permutation-invariant models) and data representations (1D grid, set, graph and complete graph).
-While some of the above studies leverage graph representations and local geometric encoding of vector polygons and state-of-art graph representation learning models to solve tasks on vector shapes, they do not yet provide a grounded, principled justification for graph-based learning methods on vector polygons.
-We attempt to fill this gap here by providing experimental results grounded in a sound theoretical motivation, and demonstrate why discrete data representations in combination with permutation-invariant learning models provide desired robustness to vector shape learning tasks.
-2.3 Graph Representation of Polygon
-Polygons are point sets connected by lines, constituting a collection of clock-wise or counter clock-wise linear rings [38].
-According to the Gestalt principles of invariance, the compact embedding of linear rings should ideally be invariant to geometric transformations (rotation, scale and shear).
-We define a graph $\mathcal{G}=(X,E)$, where $X$ is a node matrix that contains the point coordinates of geometry vertices and $E$ is an edge matrix that captures the connectivity of these vertices along the linear rings (boundaries of the outside and of any holes of polygons), hence encoding the node connectivity of the graph.
-The binary adjacency value $a_{ij}=1\in A$ is used to indicate the connectivity between nodes $i$ and $j$, with $A$ being a square adjacency matrix.
+[8] introduced ResNet1D, a 1D CNN-based polygon encoder that captures local geometric features for shape classification.
+It encodes polygon vertices using relative offsets from neighboring vertices: $\{x_{i},x_{i-1}-x_{i},x_{i+1}-x_{i},\cdots,x_{i+k}-x_{i},x_{i-k}-x_{i}\}$.
+They compared this encoding with NUFTSpec, a Non-Uniform Fourier Transform (NUFT)-based method that aggregates global shape features in the spectral domain.
+While ResNet1D focuses on shape classification, NUFTSpec is more suited for predicting topological relations.
+Their work aligns with our study in exploring different polygon encoding methods, but our focus is on evaluating transformation-invariant learning architectures across various representations (CNN, set-based, graph-based, and attention-based models).
+While previous studies explore graph-based and local geometric encoding for vector polygons, they often lack a principled justification for applying graph learning methods to polygon representations.
+We address this gap by providing: 1.
+Experimental validation grounded in theoretical motivation for robust polygon shape learning; 2.
+A principled framework demonstrating how permutation-invariant learning architectures combined with discrete data representations improve robustness to geometric transformations and vertex perturbations.
 2.3 Graph Representation Learning
@@
 By representing polygon linear rings as graphs, we can define a differentiable permutation-invariant function $\mathcal{F}$, which takes the graph representations as inputs and returns a corresponding compact graph embedding $\mathcal{F}(\mathcal{G};\theta):=\mathcal{F}(X,A;\theta)\rightarrow\mathcal{G^{\star}}\in\mathcal{R}$.
-Considering a permutation matrix $\mathcal{P}$ operating on graph $\mathcal{G}$, the permutation matrix $\mathcal{P}$ alters the order of node feature matrix $X$ in $\mathcal{G}$ along with the adjacency matrix $A$, producing a perturbed graph $\mathcal{G}^{\prime}=(PX,PAP^{\intercal})=(X^{\prime},A^{\prime})$.
+Given a permutation matrix $\mathcal{P}$ acting on a graph $\mathcal{G}$, it reorders the node feature matrix $X$ and the adjacency matrix $A$, producing a permuted graph representation $\mathcal{G}^{\prime}=(PX,PAP^{\intercal})=(X^{\prime},A^{\prime})$.
 The ideal permutation-invariant function $\mathcal{F}$ should satisfy $\mathcal{F}(X^{\prime},A^{\prime};\theta)=\mathcal{F}(X,A;\theta)$.
@@
 The local differentiable function $\phi$ linearly transforms each node feature individually $x$ to latent space and the aggregation function sums over latent node features  [^1] The basic aggregation in this setting can also be averaging or max pooling., and $\rho$ is a global differentiable function applied on the summed node features followed by a non-linear activation function.
-In this simple setting, the outputs of $\mathcal{F}$ are invariant to the permutation of node features since the aggregation sum returns the same outputs for any input permutation: $Aggr_{sum}(x_{1},x_{2},....,x_{n})=Aggr_{sum}(x_{n},...,x_{2},x_{1})$.
-This learning setting [9] makes a simple assumption on the permutations acting on input data, which can affect only individual node features without taking into account the connectivity within data (i.e., boundaries of polygon geometries).
+In this simple setting, the outputs of $\mathcal{F}$ are invariant to the permutation of node features since the summation-based aggregation returns the same outputs for any input permutation: $Aggr_{sum}(x_{1},x_{2},....,x_{n})=Aggr_{sum}(x_{n},...,x_{2},x_{1})$.
+This framework (i.e., DeepSet [9]) assumes that permutations affect only individual node features while disregarding their connectivity, which is crucial for polygonal structures.
 Graph convolution neural networks (GCN) by Kipf and Welling [16] learn convolutional features of node matrix $X$ with the normalized graph Laplacian matrix $\hat{A}=\tilde{D}^{-1/2}\tilde{A}\tilde{D}^{-1/2}$, where $\tilde{A}=A+I$ is the adjacency matrix with a identity matrix (i.e., self-connections of nodes), and $\tilde{D}$ is the diagonal degree matrix.
-Intuitively, the graph Laplacian matrix describes the divergence of energy flowing from one place (source nodes) to another (target nodes) in a graph.
-Linking this back to the linear rings of polygons and graph representation, the graph Laplacian conceptually shows similar energy divergence for nodes on the same rings(i.e., polygons with holes) and for nodes that are close to each other (i.e, polygons without holes).
+The graph Laplacian matrix describes the divergence of energy flowing from source nodes to target nodes, analogous to the spatial relationships in polygonal structures.
+Specifically, it captures connectivity patterns both within polygons (e.g., rings with holes) and between adjacent polygons.
 Thus, GCN can learn graph Laplacian embedding of node feature with a differentiable neural network layer as follows:
@@
 Here we propose to leverage the message-passing mechanism [17, 37] with the permutation-invariant function $\mathcal{F}$ to learn expressive and robust latent features of polygons.
-2.5 Message-passing Neural Network
-The principle of the message-passing mechanism is the aggregation of neighbour node features in graphs based on the “message” computed between the source and target nodes.
+3 Approach
+3.1 Graph Representation of Polygon
+Polygons are sets of points connected by lines, forming a collection of clockwise or counterclockwise linear rings [38].
+According to the Gestalt principles of invariance, the compact embedding of linear rings should ideally be invariant to geometric transformations such as rotation, scaling, and shear.
+We define a graph $\mathcal{G}=(X,E)$, where $X$ is a node matrix containing the coordinates of the geometry vertices, and $E$ is an edge matrix capturing the connectivity of these vertices along the linear rings (i.e., the boundaries of the polygon and any holes).
+The adjacency matrix $A$ encodes the connectivity between nodes, where a binary value $a_{ij}=1$ indicates that nodes $i$ and $j$ are connected.
+3.2 Message-passing Encoder
+The principle of message-passing involves aggregating neighboring node features in graphs based on messages computed between source and target nodes.
 We define the message-passing function operating on a single node as:
 $\begin{array}[]{c}h_{i}=\phi(x_{i},msg(x_{i},x_{j}),\theta),\end{array}$ (3)
-where $\{x_{i},x_{j}\}\in A=1$ and $msg(\cdot)$ is a message function with the source and target nodes as inputs.
-From Eq.3, the node feature of $x_{i}$ is updated with the computed message $(x_{i},x_{j})$, and together linearly transformed into latent space, producing the latent feature $h_{i}\in\mathcal{H}$.
-Eq. 4 present the message-passing neural network in vector view as follow:
+where $\{x_{i},x_{j}\}\in A=1$ and $msg(\cdot)$ is the message function that takes the source and target nodes as inputs.
+From Eq.3, the node feature of $x_{i}$ is updated with the computed message $(x_{i},x_{j})$ and transformed linearly into latent space, producing the latent feature $h_{i}\in\mathcal{H}$.
+The permutation-invariant message-passing encoder is represented in vector form as:
 $\begin{array}[]{c}\mathcal{F}_{msg}(X,A;\theta)=\rho_{\theta}(readout(h_{i}|h_{i}\in\mathcal{H})).\end{array}$ (4)
-A global readout function (i.e., mean or max pooling) summarizes the latent node-wise features ($\mathcal{H}$) and computes a global graph embedding for the input graph by non-linearly transforming the pooled feature.
-Hence, we propose a Messpage-passing neural network for Polygon geometries, PolyMP, as shown in Fig.
+, where a global readout function (e.g., mean or max pooling) aggregates the latent node-wise features ($\mathcal{H}$) and computes a global graph embedding by applying a non-linear transformation to the pooled features via $\rho_{\theta}$.
+Building on geometric message-passing methods for point clouds [36, 11] and polygons [35], we define a specific message-passing function $msg(\cdot)$ for PolyMP as:
+$\begin{array}[]{c}msg(x_{i},x_{j})=\alpha*\dfrac{(x_{j}-x_{i})}{std_{\mathcal{N}_{x_{i}}}+\epsilon}+\beta.\end{array}$ (5)
+, where $\alpha$ and $\beta$ are learnable parameters that adjust the local neighborhoods to a normal distribution, and $std_{\mathcal{N}_{x_{i}}}$ computes the standard deviation of the local neighborhoods $\mathcal{N}{x_{i}}$ with a small $\epsilon$ to ensure numerical stability.
+This operation facilitates stable geometric feature learning of graph neighborhoods across diverse geometric structures [39], accommodating the varying properties of polygons (Fig.5).
+After the message-passing encoder layers, a global pooling layer aggregates the node features to generate a whole-graph embedding, which is then used for downstream polygon shape classification.
+Figure 1: Model architecture of PolyMP and PolyMP-DSC.
+PolyMP consists of a simple message-passing encoder followed by a classification layer (multi-layer perceptron).
+Yellow arrows indicate the densely self-connected message-passing in PolyMP-DSC.
+Following Eq.
+3-5, we introduce a simple Message-passing neural network for Polygon geometries, named PolyMP, as shown in Fig.
 1.
-The PolyMP considers the point coordinates of polygons as input node features and the exterior connectivity of polygons as input edges.
-The message-passing layers of PolyMP then compute the “message” for each node according to the graph adjacency.
-Following existing works on point clouds [36, 11] and polygons Liu et al.
-[35], we compute the message of relative positions of two connected nodes: $msg(x_{i},x_{j})=|x_{j}-x_{i}|$, which enable aggregating the local geometric features of polygons for updating node features.
-After message-passing layers, a global pooling layer is applied over the polygons with aggregated node features to generate a whole graph embedding, which is then used for the downstream shape classification of polygons.
-Figure 1: Model architecture of PolyMP. PolyMP consists of two message-passing layer and a classification layer (Multi-layer perceptron).
-3 Experiment
-3.1 Datasets
-3.1.1 Glyph Dataset
-We introduce a synthetic dataset of highly variable geometric shapes for benchmarking the classification performance of learning models on vector polygons.
-The dataset consists of 26 Latin alphabet character glyph geometries (here also equal to semantic classes A to Z) of fonts gathered from an online source [40].
-An equivalent approach to construct a rich dataset of geometric shapes has been used to test algorithms in computational geometry before [41].
-We extract the boundaries of glyphs ( i.e., contour lines, as called in typography) for 1,413 sans serif and 1,002 serif fonts, to produce 2D simple polygon geometries compliant with OpenGeospatialConsortium [38].
-Serif and sans serif fonts are the two main typographic font families.
-Serif glyphs have salient decorative strokes ending the main character strokes increasing legibility of body text, whereas sans serif glyphs have clean strokes suitable e.g., for the display of headers.
-Importantly, these minor variations do not alter the Gestalt of the shapes (as proven by reader’s ability to consistently read the symbols).
-Thus, the labels assigned to the symbols are stable and robust.
-We encode each polygon geometry into a fixed-size feature matrix $\in R^{n\times 3}$ where each feature vector $\in R^{3}$ includes the 2D coordinates of vertices $(x,y)$ as geometric feature and a binary feature $(0,1)$ to indicate the positions of 2D coordinates, whether on outer rings or holes of polygons.
-Examples of glyph geometry are displayed in Figure 2.
-Figure 2: Glyph dataset samples. Left to right columns: A-shape to J-shape glyphs.
-3.1.2 OSM Dataset
-Latin alphabet glyph geometries share geometric similarity with building footprints.
-We use a building geometry dataset designed to test the generalizability of learning models on spatial objects proposed by Yan et al.
-[6].
-The dataset contains 10,000 real-world building footprints extracted from OSM [23], labelled by 10 categories based on a template matching to letters [42].
-The building footprints are randomly rotated and reflected from the original canonical samples (Figure 3).
-This is currently the most comprehensive real-world dataset of vector geometries with purely shape-based labels (as opposed to, e.g., building use, which should not be used for shape classification).
-We note that the assignment of labels in this dataset is somewhat subjective and likely less robust than in the glyph-based dataset.
-Note, in particular, that shape reflection is problematic for label stability, e.g., for asymmetric letters ($R$), or mirror shapes of distinct labels ($S$,$Z$).
-Figure 3: OSM dataset samples.
-Top row: building geometries of standard shapes.
-Bottom row: geometric transformation counterparts.
-Left to right columns: E, F, H, I, L, O, T, U, Y and Z-shape buildings.
-[6]
-3.1.3 Pre-processing
-Label-preserving transformations
-To examine the invariances of geometric transformations of polygon representations and corresponding learning models with respects to data transformations, we consider four training datasets under data transformations, each consisting of 20%, 40%, 60% and 80% of glyph polygons that have undergone label-preserving geometric transformations.
-The details of transformations are presented in Table 1.
-The original and transformed data samples are visualised in Fig.
-4.
-We thus obtain five training datasets, including the original dataset: dataset 0%, 20%, 40%, 60% and 80% for the experiment.
-The test set consists of data randomly sampled from the five datasets.
-Table 1: Label-preserving transformations applied to samples of Glyph dataset maintains the semantic information (label) of original polygons (i.e., 180 degrees rotation of letter M alter its semantic label, converting to letter W).
-Transformation Operation
-Rotation Rotate data around its centroid by a random angle $\in$ $\left[-75^{\circ},75^{\circ}\right]$.
-Scaling Scale data by random distinct factors $\in$ $\left[0.1,2\right]$ on the $x$ and $y$ axes.
-Shearing Shear data in a random angle $\in$ $\left[-45^{\circ},45^{\circ}\right]$ on the $x$ and $y$ axes.
-Figure 4: Data samples of Glyph dataset under label-preserving transformations.
-Polygon Simplifications
-To test the invariances to trivial vertices of polygon, we apply the Douglas–Peucker algorithm [43] (tolerance = 1.0) on the original size of glyph polygon (in the order of 50 $\times$ 50 units) to generate simplified samples for glyph dataset.
-The resulting polygon linestrings are defined by fewer vertices (in particular, excluding all co-linear or close to co-linear vertices).
-Notice that the simplification occurs before the normalisation to the range of (-1,1) and any augmentation of polygons, ensuring that the simplified geometries preserve the topology and essence of the shape of the original glyph (see Fig.
-5).
-Figure 5: Visualisation of original and simplified polygon examples generated with Douglas–Peucker algorithm [43].
-The simplified polygons, compared to original polygons, contain a subset of the vertices that defined the original curve.
-3.2 Training Configuration
-We evaluate the performances of learning models from previous studies, in particular: the VeerCNN in Veer et al.
-[2], based on 1D sequences converted from polygons as inputs; the DeepSet from Zaheer et al.
-[9] and SetTransformer Lee et al.
-[12], based on polygon vertices as input point clouds; and the GCAE in Yan et al.
-[6], representing polygons as input graphs.
-We then compare their performances with the proposed message-passing model for polygons (PolyMP).
-The model configurations used in our experiments are in Table 2.
-Correspondingly, we depict the data representations of polygons and baseline models in Figure 6.
-The proposed PolyMP is demonstrated in Figure 1.
-For experimental control, every model tested was constrained to a comparable model depth and latent space of feature embedding.
-For model training on Glyph dataset and fine-tuning on OSM dataset, we compute the cross entropy loss [44] to determine training gradients for fast training convergence and good model generalizability.
-We use the Adam optimization [45] with an initial learning rate of $0.01$.
-For better model generalization, we adopt a learning rate decay strategy that reduces the learning rate by a factor of $10$ once the model training reaches a performance plateau over $25$ epochs.
-An early-stopping mechanism halts the training once the loss ceases to decrease for $50$ epochs, to mitigate overfitting.
-We set the batch size to $64$ and the total training epochs to $100$.
-Thus, importantly, the models here are not designed to individually outperform current state-of-the-art for each architecture, but to maintain comparability across architectures and to enable evaluation of the effects of data representations on the results.
+PolyMP processes polygon point coordinates as input node features and the connectivity of polygon edges as input edges.
+The message-passing encoders of PolyMP compute the “message” for each node based on the graph adjacency and update the node features in each layer.
+3.3 Densely Self-Connected Message-Passing
+As noted in recent studies of graph message-passing networks [40], the basic message-passing mechanism restricts information flow due to the absence of self-loop information in graph representation learning.
+To address this limitation and enhance PolyMP’s learning capability, we adopt a densely self-connected (DSC) message-passing mechanism [41, 40, 18], defined as:
+$\begin{array}[]{c}\mathcal{G}_{l}=f(\rho_{\theta}(readout(h_{i}|h_{i}\in\mathcal{H}_{l}))+\mathcal{G}_{l-1}+\mathcal{G}_{l-2}+...+\mathcal{G}_{0}),\end{array}$ (6)
+where we extend the basic PolyMP model to PolyMP-DSC, as shown in Fig.
+1.
+The PolyMP-DSC aggregates latent graph-wise features $\mathcal{G}_{l}$ at layer $l$ by incorporating latent graph-wise features $\mathcal{G}_{l-1...0}$ from previous message-passing layers.
+This enables the model to capture hierarchical self-looped graph information, resulting in richer graph-wise representations.
+3.4 Implementation
+Figure 2: Data representations of polygons (as 1D sequence, Point Set, Graph and Complete Graph) and the corresponding learning models (Baselines: VeerCNN, DeepSet, GCAE and SetTransformer).
 Table 1: Model configuration.
@@
 GCAE [6] 2 {2, 64, 64} 7,613 Global avg.
-PolyMP (Ours) 2 {2, 64, 64} 11,837 Global avg.
-Figure 6: Data representations of polygons (as 1D sequence, Point Set, Graph and Complete Graph) and the corresponding learning models (Baselines: VeerCNN, DeepSet, GCAE and SetTransformer).
+NUFTSpec [8] 2 {288, 128, 64} 48,635 -
+DSC-NMP [40] 2 {2, 64, 64} 20,539 Global add.
+PolyMP (Ours) 2 {2, 64, 64} 11,837 Global max
+PolyMP-DSC (Ours) 2 {2, 64, 64} 16,579 Global max
+We evaluate the performance of learning models from previous studies, specifically: the VeerCNN model from Veer et al.
+[2], which uses 1D sequences converted from polygons as input; the DeepSet model from Zaheer et al.
+[9] and SetTransformer model from Lee et al.
+[12], which treat polygon vertices as input point clouds; the GCAE model from Yan et al.
+[6], which represents polygons as input graphs; and the DSC-NMP model from Fan et al.
+[40], a baseline graph learning model that processes polygons as input graphs and applies DSC neural message-passing for graph feature learning.
+We also evaluate the NUFTSpec model from Mai et al.
+[8], which encodes polygon geometries by converting them into feature vectors in the rasterized spectral domain.
+These models are compared against the proposed graph message-passing models for polygons, PolyMP and PolyMP-DSC.
+The model configurations used in our experiments are summarized in Table 1, and the data representations of polygons and baseline learning models are illustrated in Figure 2.
+For experimental control, every model trained and tested was constrained to a comparable model depth and latent space of feature embedding.
+During model training on the Glyph dataset and fine-tuning on the OSM dataset, we compute cross-entropy loss [42] to determine training gradients for fast convergence and good model generalizability.
+We use Adam optimization [43] with an initial learning rate of 0.01.
+For better model generalization, we adopt a learning rate decay strategy that reduces the learning rate by a factor of 10 once the model training reaches a performance plateau after 25 epochs.
+An early-stopping mechanism halts the training once the loss ceases to decrease for 50 epochs, mitigating overfitting.
+The batch size is set to 64, and the total training epochs are set to 100.
+Importantly, the models are not designed to individually outperform state-of-the-art models for each architecture, but rather to maintain comparability across architectures and enable the evaluation of the effects of data representations on results.
+4 Experiment
+4.1 Glyph Dataset
+We introduce a synthetic dataset of highly variable geometric shapes to benchmark the classification performance of learning models on vector polygons.
+This dataset consists of 26 Latin alphabet character glyph geometries, representing semantic classes from A to Z, gathered from an online source [44].
+A similar approach has been used previously in computational geometry to construct rich datasets for algorithm testing [45].
+The synthetic dataset includes the boundaries (contour lines) of glyphs extracted from 1,413 sans serif and 1,002 serif fonts, resulting in 2D simple polygon geometries compliant with the standards of OpenGeospatialConsortium [38].
+Serif and sans serif fonts are the two primary typographic families.
+Serif glyphs feature decorative strokes that enhance the legibility of body text, while sans serif glyphs have clean, minimal strokes, making them more suitable for headers.
+Importantly, these minor variations do not affect the overall Gestalt of the shapes, as readers can consistently recognize the symbols.
+This stability ensures that the labels assigned to the symbols are both stable and robust.
+Each polygon geometry is encoded into a fixed-size feature matrix of size $\in\mathbb{R}^{n\times 3}$, where each feature vector $\in\mathbb{R}^{3}$ contains the 2D coordinates of vertices $(x,y)$ as the geometric feature, and a binary feature $(0,1)$ to indicate the position of each 2D coordinate (whether it lies on the outer rings or holes of the polygons).
+Examples of glyph geometries are shown in Figure 3.
+Figure 3: Glyph dataset samples. Left to right columns: A-shape to J-shape glyphs.
+4.2 OSM Dataset
+Latin alphabet glyph geometries share geometric similarities with building footprints.
+To test the generalizability of learning models on spatial objects, we use a building geometry dataset proposed by Yan et al.
+[6].
+The dataset contains 10,000 real-world building footprints extracted from OpenStreetMap (OSM) [23], labeled into 10 categories based on template matching to letters [46].
+The building footprints are randomly rotated and reflected from the original canonical samples, as shown in Figure 4.
+This dataset is currently the most comprehensive real-world collection of vector geometries with purely shape-based labels (as opposed to labels based on building use, which should not be used for shape classification).
+However, we note that the assignment of labels in this dataset is somewhat subjective and may be less robust compared to the glyph-based dataset.
+Specifically, shape reflection can pose challenges for label stability, especially for asymmetric letters (e.g., $R$) or mirror image shapes of distinct labels (e.g., $S$ and $Z$).
+Figure 4: OSM dataset samples.
+Top row: building geometries of standard shapes.
+Bottom row: geometric transformation counterparts.
+Left to right columns: E, F, H, I, L, O, T, U, Y and Z-shape buildings.
+[6]
+4.3 Pre-processing
+Label-preserving transformations
+To assess the invariance of geometric transformations in polygon representations and their effect on learning models, we apply label-preserving geometric transformations to the glyph polygons.
+We create four transformed training datasets, each consisting of 20%, 40%, 60%, and 80% of glyph polygons that have undergone these transformations.
+The details of the transformations are listed in Table 2, with the original and transformed data samples visualized in Fig.
+5.
+This results in five training datasets: 0%, 20%, 40%, 60%, and 80% transformations, which will be used for the experiment.
+The test set consists of data randomly sampled from these five datasets.
+Table 2: Label-preserving transformations applied to samples of Glyph dataset maintains the semantic information (label) of original polygons (i.e., 180 degrees rotation of letter M alter its semantic label, converting to letter W).
+Transformation Operation
+Rotation Rotate data around its centroid by a random angle $\in$ $\left[-75^{\circ},75^{\circ}\right]$.
+Scaling Scale data by random distinct factors $\in$ $\left[0.1,2\right]$ on the $x$ and $y$ axes.
+Shearing Shear data in a random angle $\in$ $\left[-45^{\circ},45^{\circ}\right]$ on the $x$ and $y$ axes.
+Figure 5: Data samples of Glyph dataset under label-preserving transformations.
+Polygon Simplifications
+To assess the invariance of polygon representations to trivial vertices, we apply the Douglas–Peucker algorithm [47] with a tolerance of 1.0 to the original glyph polygons (typically sized around 50 $\times$ 50 units).
+This simplification process generates polygon samples with fewer vertices, particularly by excluding co-linear or nearly co-linear vertices.
+Importantly, the simplification is applied before normalizing the polygons to the range of (-1, 1) and before any augmentation, ensuring that the simplified geometries preserve the topology and essential shape of the original glyph (see Fig.
+6).
+Figure 6: Visualisation of original and simplified polygon examples generated with Douglas–Peucker algorithm [47].
+The simplified polygons, compared to original polygons, contain a subset of the vertices that defined the original curve.
 5 Result
@@
 5.1 Results on Glyph Dataset
-Table 3 and 4 demonstrates the model performances testing on Glyph dataset and the simplified polygons, respectively.
-At $0\%$ transformation ratio in Table 3, we observe that the PolyMP achieves $99.41\%$, $42.91\%$ and $80.03\%$ acc.
-on the original, rotated and scaled test samples, whereas VeerCNN achieves $59.07\%$ acc.
-on sheared samples.
-While deepset, set transformer and GCAE maintain comparable performances ($94.62\%$, $65.01\%$; and $96.25\%$, $67.91\%$; and $95.16\%$, $64.54\%$, respectively) on original and scaled test samples, the models suffer large performance deterioration, achieving only $23.67\%$, $26.00\%$ and $25.23\%$, $27.79\%$ and $31.38\%$ and $38.84\%$ test acc.
-on rotated and sheared samples, respectively.
-From the overall accuracy, the VeerCNN outperforms the PolyPM by a slight margin ($69.43\%$ and $69.01\%$).
-At 20% transformation ratio, the PolyMP improves model test acc.
-on rotated, scaled and sheared samples rapidly, achieving 83.17% (+40.26%), 91.29% (+11.26%) and 86.86% (+33.54%) respectively.
-In overall acc., the PolyMP achieves 90.12% O.A.
-(+21.11%) and outperforms VeerCNN.
-In comparison, the other models can not match with PolyMP and achieve comparable performances from increasing data diversity of training set.
-At 40% and 60% transformation ratio, we notice that the accuracy improvements gained from increasing the ratio of geometric transformed samples in training set have converged.
-In this case, the PolyMP maintains its leading performances among models (91.91% and 93.06% O.A.), and VeerCNN (83.30% and 84.88% O.A.), deepset (80.12% and 81.49% O.A.), set transformer (80.33% and 81.94% O.A.) and GCAE (72.72% and 75.50% O.A.) fall behind.
-At 80% transformation ratio, the PolyMP achieves the optimal test performances at rotated and sheared samples, giving 90.35% and 91.52% test accuracy.
-This empirical results suggest that compared to the sequence encoding of polygons and CNN as well as the set representation and set-based learning models, the graph representation together with message-passing neural network (i.e., PolyMP) are robust to the permutations of feed-in order of polygon vertices, which occurs in geometric transmutations of polygons, in particular rotation and shearing.
-From a different perspective, the ratio of original samples is decreasing in the training set progressively from the 0% to 80% transformation ratios datasets.
-The PolyMP only encounters 1.46% test performance deterioration along this progression (99.41% to 97.95% acc.), while the VeerCNN (-4.13% acc.), deepset (-3.96% acc.), set transformer (-5.08% acc.) and GCAE (-5.58% acc.) suffer larger performance drops on the test accuracy of original samples.
 Table 3: Test performances on Glyph dataset.
@@
 GCAE 95.16 31.38 64.54 38.84 57.83
-PolyMP 99.41 42.91 80.03 53.32 69.01
+NUFTSpec 98.42 34.90 72.07 39.68 61.34
+DSC-NMP 96.81 35.71 72.83 43.21 62.24
+PolyMP 99.68 39.91 82.35 53.22 69.03
+PolyMP-DSC 99.58 41.74 86.45 56.16 71.21
 20% VeerCNN 95.12 67.63 83.53 77.78 81.05
@@
 GCAE 93.31 47.31 71.47 57.11 67.61
-PolyMP 98.92 83.17 91.29 86.86 90.12
+NUFTSpec 98.00 65.27 84.03 70.74 79.61
+DSC-NMP 93.68 52.38 78.86 61.72 71.73
+PolyMP 99.10 84.89 93.26 88.17 91.30
+PolyMP-DSC 99.19 86.20 94.60 90.65 92.64
 40% VeerCNN 94.75 73.26 84.77 80.29 83.30
@@
 GCAE 91.70 59.55 74.35 64.42 72.72
-PolyMP 98.48 87.06 92.49 89.25 91.91
+NUFTSpec 97.44 78.84 86.00 78.71 85.24
+DSC-NMP 86.71 54.01 71.96 57.97 67.71
+PolyMP 98.85 88.53 94.35 90.74 93.11
+PolyMP-DSC 99.05 90.18 95.33 92.57 94.18
 60% VeerCNN 94.50 77.16 85.38 82.36 84.88
@@
 GCAE 90.46 65.02 76.38 68.78 75.50
-PolyMP 98.30 89.40 93.43 91.11 93.06
+NUFTSpec 96.80 83.13 88.40 81.76 87.55
+DSC-NMP 90.97 64.17 80.00 67.73 75.76
+PolyMP 98.59 90.08 94.11 91.65 93.69
+PolyMP-DSC 98.79 91.80 95.76 93.20 94.91
 80% VeerCNN 93.51 78.97 85.25 83.17 85.25
@@
 GCAE 89.58 67.95 77.99 70.55 76.83
-PolyMP 97.95 90.35 93.30 91.52 93.42
-In Table 4, the test performances of models trained and evaluated on the simplified polygons of Glyph dataset are presented.
-In this experiment setting, we focus on models’ test performances on simplified polygons containing fewer numbers of vertices, and compare them with the test performances showed in Table 3.
-We observe similar test performance trends to those of models evaluated on the simplified polygons.
-Noticeably, there is a significant performance deterioration in VeerCNN compared with the test performances in Table 3 on simplified polygons.
-Specifically, at 0% transformation ratio, the test performances of VeerCNN drops from 69.43% O.A.
-(Table 3) to 66.56% O.A.
-(Table 4), i.e., a 2.87% performance deterioration.
-As the data diversity increases, the performance deterioration in VeerCNN increases, too, from 5.2%, progressively through to 5.56% O.A.
-deterioration respectively from 20% to 80% transformation ratios.
-The performance deterioration of VeerCNN strongly indicates that CNN models are not robust to varying the numbers of vertices in polygons.
-This can be attributed to a number of factors, incl.
-the data encoding of polygons in VeerCNN as as fixed-length 1D sequences with zero padding.
-Simplifying polygons and removing trivial vertices from polygons alters the length of 1D sequences, and hence the geometric information of simplified polygons are not only diluted by the zero padding in the 1D sequences, but also by alteration of vertex adjacency explicitly encoded in the 1D sequence.
-In comparison, we notice only small performance drops occur in general set-based and graph-based learning models.
-We credit this to set and graph representation of polygons and the corresponding learning models as both set and graph-based models accept input polygons of varying lengths.
-In particular, the graph-based message-passing network (i.e., PolyMP) learns salient local node features from connected neighbours in the message-passing function with permutation-invariant pooling operations (i.e., Max-pooling), as shown in E.q.
-4.
-A similar pooling operation is also adopted in set-based learning models [9, 12].
+NUFTSpec 96.15 84.40 89.24 83.46 88.32
+DSC-NMP 94.31 74.31 86.75 78.36 83.46
+PolyMP 98.48 91.64 94.69 92.36 94.31
+PolyMP-DSC 98.93 93.19 95.95 94.26 95.61
 Table 4: Test performances on Glyph dataset, consisting of simplified polygons generated by Douglas–Peucker algorithm (tolerant=1.0) with the removal of trivial vertices.
@@
 GCAE 94.83 31.64 64.21 39.90 58.08
-PolyMP 98.89 43.71 79.15 52.70 68.63
+NUFTSpec 98.03 34.91 71.96 40.07 61.30
+DSC-NMP 95.40 35.21 71.61 41.56 61.04
+PolyMP 99.27 40.52 81.90 53.50 68.92
+PolyMP-DSC 99.36 42.26 86.16 56.68 71.07
 20% VeerCNN 90.55 62.58 77.80 72.31 75.85
@@
 GCAE 92.26 47.29 71.28 57.33 67.39
-PolyMP 97.61 82.45 88.42 84.84 88.22
+NUFTSpec 97.76 65.10 83.97 71.41 79.65
+DSC-NMP 92.20 51.73 77.60 60.44 70.55
+PolyMP 98.37 83.37 91.41 87.25 89.74
+PolyMP-DSC 98.93 84.57 93.56 89.15 91.19
 40% VeerCNN 89.63 67.82 78.72 74.61 77.73
@@
 GCAE 91.65 59.30 74.95 64.79 72.94
-PolyMP 97.55 86.14 89.99 87.29 90.21
+NUFTSpec 97.23 78.85 86.09 78.95 85.25
+DSC-NMP 85.08 52.92 69.90 56.73 66.21
+PolyMP 98.10 86.47 92.72 88.73 91.09
+PolyMP-DSC 98.66 89.04 94.45 91.84 93.28
 60% VeerCNN 89.29 71.91 79.42 77.29 79.50
@@
 GCAE 90.06 63.61 76.64 68.05 74.79
-PolyMP 97.35 88.73 91.04 89.31 91.47
+NUFTSpec 96.62 82.49 88.48 82.16 87.46
+DSC-NMP 90.02 62.99 79.47 66.67 74.83
+PolyMP 97.74 89.50 92.99 90.59 92.48
+PolyMP-DSC 98.38 91.53 94.89 92.76 94.13
 80% VeerCNN 88.05 73.89 79.55 77.17 79.69
@@
 GCAE 88.22 66.55 78.01 70.44 76.06
-PolyMP 97.01 90.19 91.53 90.31 92.22
+NUFTSpec 96.10 84.56 89.16 83.28 88.27
+DSC-NMP 93.06 73.10 85.03 76.99 82.08
+PolyMP 97.42 90.55 93.25 91.36 92.97
+PolyMP-DSC 98.47 92.78 95.10 93.43 94.77
+Table 3 and Table 4 present the model performances on the Glyph dataset and simplified polygons, respectively.
+At the $0\%$ transformation ratio in Table 3, we observe that PolyMP achieves $99.68\%$, $39.91\%$, and $82.35\%$ accuracy on the original, rotated, and scaled test samples, respectively, while VeerCNN achieves $59.07\%$ accuracy on sheared samples.
+In comparison, DeepSet, SetTransformer, and GCAE maintain comparable performance on the original and scaled test samples ($94.62\%$, $65.01\%$; $96.25\%$, $67.91\%$; and $95.16\%$, $64.54\%$, respectively).
+However, models that learn from set (DeepSet and SetTransformer) and graph inputs (GCAE and DSC-NMP) experience significant performance deterioration on rotated and sheared samples, with accuracy of $23.67\%$, $26.00\%$; $25.23\%$, $27.79\%$; $31.38\%$, $38.84\%$; and $35.71\%$, $43.21\%$, respectively.
+Overall, PolyMP-DSC, which learns densely connected graph features, achieves the highest classification accuracy at $71.21\%$, significantly outperforming the baseline DSC-NMP model ($62.24\%$ O.A.).
+At the 20% transformation ratio, both PolyMP (84.89% (+44.98%), 93.26% (+10.91%), and 88.17% (+34.95%)) and PolyMP-DSC (86.20% (+44.46%), 94.60% (+8.15%), and 90.65% (+34.49%)) show significant improvements in test accuracy on rotated, scaled, and sheared samples, respectively.
+In terms of overall accuracy, PolyMP achieves 91.30% O.A.
+(+22.27%), outperforming VeerCNN’s 81.05% O.A.
+PolyMP-DSC records the highest O.A.
+at 92.64% (+21.43%).
+In contrast, the baseline learning models are unable to match the performance of PolyMP and PolyMP-DSC, despite the increase in training set diversity.
+At the 40% and 60% transformation ratios, we observe that the accuracy improvements from increasing the proportion of geometrically transformed samples in the training set have plateaued.
+In this case, PolyMP achieves 93.11% and 93.69% O.A., while PolyMP-DSC leads with 94.18% and 94.91% O.A.
+Both models maintain their superior performance compared to others.
+VeerCNN records 83.30% and 84.88% O.A., followed by DeepSet (80.12% and 81.49% O.A.), SetTransformer (80.33% and 81.94% O.A.), GCAE (72.72% and 75.50% O.A.), DSC-NMP (57.97% and 67.73% O.A.), and NUFTSpec (70.74% and 78.71% O.A.), all of which fall behind.
+At the 80% transformation ratio, PolyMP and PolyMP-DSC achieve the highest test performances on rotated and sheared samples, with accuracy of 91.64%, 92.36%, and 93.19%, 94.26%, respectively.
+These empirical results suggest that, compared to sequence encoding of polygons using CNNs and set representations with set-based learning models, graph representations combined with message-passing neural networks (PolyMP and PolyMP-DSC) are more robust to permutations of polygon vertices.
+This robustness is particularly evident in geometric transformations such as rotation and shearing, where the feed-in order of vertices can vary.
+Comparing basic PolyMP with the extended PolyMP-DSC, we observe testing performance improvements of up to $2\%$ O.A.
+on the Glyph dataset.
+These results highlight the effectiveness of incorporating hierarchical self-looped graph representations via densely self-connected message passing for polygon encoding and shape-based classification.
+From a different perspective, as the proportion of original samples progressively decreases in the training set (from 0% to 80% transformation ratios), PolyMP and PolyMP-DSC experience only a slight test performance deterioration, with a decrease from 99.68% to 98.48% (-1.20%) and 99.58% to 98.93% (-0.65%), respectively.
+In contrast, VeerCNN (-4.13%), DeepSet (-3.96%), SetTransformer (-5.08%), GCAE (-5.58%), NUFTSpec (-2.27%), and DSC-NMP (-2.50%) exhibit more significant performance drops in test accuracy on the original samples.
+In Table 4, we present the test performances of models trained and evaluated on simplified polygons from the Glyph dataset.
+This experimental setting focuses on evaluating the models’ performance on polygons with a reduced number of trivial vertices (Fig.
+6) and compares these results with those shown in Table 3.
+We observe similar performance trends in the models when evaluated on the simplified polygons.
+Notably, VeerCNN experiences a significant drop in performance compared to the results in Table 3.
+Specifically, at the 0% transformation ratio, VeerCNN’s test performance decreases from 69.43% O.A.
+(Table 3) to 66.56% O.A.
+(Table 4), representing a 2.87% decline.
+As data diversity increases, the performance deterioration in VeerCNN becomes more pronounced, with a drop of 5.2%, progressively increasing to 5.56% O.A.
+at transformation ratios from 20% to 80%.
+The performance deterioration of VeerCNN strongly suggests that CNN models are not robust to changes in the number of vertices in polygons.
+This can be attributed to several factors, including the data encoding method used in VeerCNN, which represents polygons as fixed-length 1D sequences with zero padding.
+Simplifying polygons by removing trivial vertices alters the length of these sequences, which not only dilutes the geometric information through zero padding but also disrupts the explicit encoding of vertex adjacency within the 1D sequences.
+In comparison, we observe only minor performance drops in general set-based and graph-based learning models.
+This can be attributed to the flexibility of set and graph representations of polygons, which allow these models to accept input polygons of varying lengths.
+Specifically, graph-based message-passing networks (i.e., PolyMP and PolyMP-DSC) effectively learn local geometric features from neighboring nodes using a message-passing mechanism, followed by permutation-invariant pooling operations (e.g., max-pooling), as shown in Eq.
+4.
+A similar pooling operation is utilized in set-based learning models [9, 12].
 5.2 Results on OSM Dataset
-We fine-tune models pre-trained on the glyph 80% dataset with the OSM dataset, and report the classification test accuracy of fine-tuned models in Table 5 for building footprints with original and normalised coordinates respectively.
-PolyMP achieves optimal test performance on normalised footprints with $91.92\%$ O.A.
-on the OSM dataset, and performs better than all other methods.
-Both set-based methods (i.e., DeepSet: $95.71\%$ and SetTransformer: $94.99\%$) and graph-based GCAE ($94.79\%$) achieve comparable performance on OSM-O dataset with building footprints without geometric transformations.
-Conversely, the convolutional VeerCNN shows lower performance of $90.54\%$.
-When testing on OSM-R with randomly rotated and reflected footprints, PolyMP maintains a robust performance of $85.62\%$ and the baseline methods suffer performance deterioration at various levels (VeerCNN: $90.54\%\rightarrow 78.51\%$, DeepSet: $95.71\%\rightarrow 84.86\%$, SetTransformer: $94.99\%\rightarrow 83.87\%$, GCAE: $94.79\%\rightarrow 80.19\%$).
-Performances of models fine-tuned on OSM footprints with original spatial coordinates drop for all models except for PolyMP, which maintains a performance of $81.58\%$ O.A..
-This robust performance of PolyMP against the translation of spatial coordinates can be accounted for by the message-passing network of PolyMP where the local geometric features of polygons are computed (i.e., the relative positions of polygons $msg(x_{i},x_{j})=|x_{j}-x_{i}|$).
-This relative positional features reduce the impact of absolute positioning and effectively normalise the spatial coordinates of polygons while preserving the geometric properties of polygons.
-VeerCNN outperforms DeepSet ($24.34\%$ O.A) and CGAE ($23.18\%$ O.A) and achieves a test performance of $42.68\%$ O.A.
-when fine-tuned on the original OSM footprints.
-We account for VeerCNN’s performance by the built-in translation invariant priors in conventional convolution neural networks.
-Similarly, SetTransformer shows comparable performance as VeerCNN ($44.72\%$ O.A), which can be attributed to the self-attention features based on the fully connected graph of point sets.
+Table 5: Fine-tuned test performance on the OSM dataset.
+OSM-O and OSM-R indicate the classification accuracy of models tested on original and rotated & reflected samples, respectively.
+Overall accuracy is reported as OSM O.A.
+Model OSM-O Acc. OSM-R Acc. OSM O.A.
+VeerCNN 86.30 72.44 79.36 Normalised coordinates $\in(-1,1)$
+DeepSet 94.79 82.31 88.62
+SetTransformer 89.10 68.73 78.90
+GCAE 95.23 78.15 86.94
+NUFTSpec 88.02 60.66 74.26
+DSC-NMP 87.58 66.73 77.14
+PolyMP 96.59 81.03 88.58
+PolyMP-DSC 95.75 78.63 87.20
+VeerCNN 42.23 42.09 42.16 Original coordinates
+DeepSet 22.00 21.65 21.84
+SetTransformer 35.62 35.30 35.46
+GCAE 22.12 22.32 22.22
+NUFTSpec 88.02 61.38 74.42
+DSC-NMP 31.81 33.23 32.52
+PolyMP 50.64 39.86 46.20
+PolyMP-DSC 44.19 36.78 40.68
+The results in Table 5 show the test accuracy of fine-tuned classifiers on real-world building footprints from the OSM dataset, using feature encoders pre-trained on the Glyph dataset (80% subset).
+These fine-tuned models were evaluated on building footprints with both normalized and original coordinates.
+PolyMP achieves an accuracy of 88.58% for building footprints with normalized coordinates and 46.20% for footprints with original coordinates.
+PolyMP-DSC demonstrates slightly lower test accuracy, with 87.20% for normalized coordinates and 40.68% for original coordinates, but still outperforms other baseline methods, except for DeepSet.
+This can be attributed to the ability of both models to compute local geometric features via message-passing networks, reducing the impact of absolute positioning while preserving polygon geometry.
+DeepSet performs best, with 88.62% O.A.
+on normalized coordinates and 82.31% O.A.
+on rotated and reflected buildings, making it the top performer in some cases.
+The NUFTSpec model, which transforms polygon geometries into consistent spectral feature vectors via Non-Uniform Fourier Transforms, demonstrates stable performance at approximately 74% O.A.
+across both normalized and original coordinates.
+This stability stems from a series of affine transformations (e.g., scaling and translation) applied to input polygons during feature transformation, ensuring that each polygon is projected into a unit space and positioned within the same relative coordinate system, as discussed in [8].
 In Fig.
-7, we visualise the classification results of PolyMP evaluated on spatial building footprints in OSM-O (Fig.
-7 (a)) and OSM-R (Fig.
-7 (b)).
-Predictions indicate that the fine-tuned PolyMP can robustly classify building footprints on a map in varying sizes and geometric shapes based on the geometric invariant features learnt by message-passing networks, taking the graph-based representation of polygons as inputs.
-Table 5: Fine-tune test performances on OSM dataset.
-OSM-O and OSM-R acc.
-suggest the the classification accuracy of models tested on original and rotated & reflected samples.
-Overall accuracy is reported for OSM O.A..
-Model OSM-O Acc. OSM-R Acc. OSM O.A.
-VeerCNN 90.54 78.51 84.52 Normalised coordinates $\in(-1,1)$
-DeepSet 95.71 84.86 90.56
-SetTransformer 94.99 83.87 89.42
-GCAE 94.79 80.19 88.02
-PolyMP 98.04 85.62 91.92
-VeerCNN 42.35 43.01 42.68 Original coordinates
-DeepSet 24.12 24.08 24.34
-SetTransformer 44.67 44.77 44.72
-GCAE 23.44 23.04 23.18
-PolyMP 89.86 73.56 81.58
+7, classification results of PolyMP on building footprints from the two OSM subsets (OSM-O and OSM-R) are visualized.
+The results show PolyMP’s robustness in classifying building footprints of varying sizes and geometric shapes, leveraging the geometric invariant features learned through message-passing networks.
+These findings underline the effectiveness of graph message-passing models, particularly PolyMP and PolyMP-DSC, in handling both normalized and original coordinate systems and achieving strong performance across various data transformations.
 (a) OSM-O
 (b) OSM-R
-Figure 7: Shape classification of building footprints by PolyMP on the selected region in OSM dataset (downtown areas of Shanghai, China [6]). Model predictions are colored in red.
+Figure 7: Shape classification of building footprints using PolyMP on a selected region of the OSM dataset [6]. Model predictions are highlighted in red.
 6 Discussion
-(a) Glyph dataset.
-(b) OSM dataset.
-Figure 8: Failure cases of PolyMP on Glyph and OSM datasets.
 (a) Sheared M glyph.
@@
 (f) Original Q glyph.
-Figure 9: Feature maps to visualize salient features learned by individual models to classify input geometries to output categories.
+Figure 8: Feature maps to visualize geometric features learned by individual models to classify input geometries to output categories.
 The darker the color, the more salient are the features learned.
 Feature values are normalised $\in[0,1]$.
+(a) Glyph dataset.
+(b) OSM dataset.
+Figure 9: Failure cases of PolyMP on Glyph and OSM datasets.
 6.1 Model Expressivity
-For the interpretation of model expressivity with respect to various polygon encodings and learning models, we visualize the feature maps of models tested on original and geometric transformed samples in glyph dataset in Figure 9.
-Figs.
-9a and 9b depict the salient feature maps and class predictions of four models (Deepset, Set Transformer, GCAE and PolyMP) on geometric transformed samples (sheared and rotated glyph M).
-In this case, the PolyMP correctly predicts the class label of glyph $M$, whereas DeepSet, Set Transformer and GCAE have made incorrect predictions on the geometric transformed samples.
-By comparing the feature maps in Figures 9c and 9d, the PolyMP learns to identify the global structures or the “skeleton” of polygons via non-trivial vertices on the boundary.
-The message-passing network and local pooling function (i.e., max-pooling) can capture the salient vertices from local neighbours (i.e., non-trivial vertices (dark color) in the clusters of trivial vertices (light color), which enable a robust embedding of polygons under geometric transformations.
-Set-based models and GCAE learn sparser geometric features for input polygons, capturing salient features regionally in node clusters and ignoring geometric features of more distant nodes.
-This can be attributed to the set representation where any relational information between nodes is neglected.
-Figures 9e and 9f demonstrate the feature maps on polygons with holes.
-Both Deepset and Set Transformer only capture geometric features from non-trivial vertices located at the exterior boundary of polygons.
-While set-based models accept point set encoding of polygons in varying length and the output features are invariant to the feed-in order of vertices, such methods largely ignore the connectives and interactions between points.
-Such a learning setting inherently hinders the model’s capability to learn geometric information from polygons with holes.
-The graph convolution model, GCAE, lacks learning flexibility in learning local geometric features as discussed in the previous sections.
-On the other hand, PolyMP can capture geometric features from salient points located at the exterior and interiors of polygons.
-In the message-passing network, the exterior and interior of polygons are considered to be two distinct elements as there is no direct connection between points located at the two boundaries.
-In other words, the geometric features (i.e., the “message”) are only passed between points that are connected in a connected component of a graph, in this case defined by the linear rings of polygon boundaries.
+To interpret the expressivity of different polygon encoding and learning models, we visualize the feature maps of models tested on original and geometrically transformed samples from the Glyph dataset in Figure 8.
+Figures 8a and 8b illustrate the salient feature maps and class predictions of set-based (DeepSet, Set Transformer) and graph-based (GCAE, DSC-NMP, PolyMP, and PolyMP-DSC) learning models on geometrically transformed samples.
+Notably, PolyMP and PolyMP-DSC correctly predict the class label of the sheared and rotated glyph $M$, whereas DeepSet, Set Transformer, GCAE, and DSC-NMP make incorrect predictions.
+By comparing the feature maps in Figures 8c and 8d, we observe that PolyMP and PolyMP-DSC effectively capture the global structure or “skeleton” of polygons by identifying salient, non-trivial vertices along the boundary.
+The message-passing encoder, coupled with a local pooling function (i.e., max-pooling), extracts geometric features from local neighborhoods, where non-trivial vertices (darker color) are distinguished within clusters of trivial vertices (lighter color).
+This enables the models to generate robust embeddings of polygons, even under geometric transformations.
+In contrast, set-based models and GCAE learn sparser geometric representations, capturing regional features in node clusters while neglecting relational information between more distant nodes.
+This limitation arises from their reliance on set representations, which inherently discard connectivity information between vertices.
+Figures 8e and 8f illustrate the feature maps for polygons with holes.
+DeepSet and Set Transformer primarily capture geometric features from non-trivial vertices along the exterior boundary, while largely ignoring those within the interior.
+Although set-based models can encode polygons of varying lengths and produce permutation-invariant feature representations, their lack of connectivity modeling restricts their ability to learn geometric structures in polygons with holes.
+Similarly, the graph convolution model GCAE struggles to flexibly capture local geometric features, as discussed in previous sections.
+In contrast, PolyMP and PolyMP-DSC effectively capture geometric features from both the exterior and interior boundaries of polygons.
+This is achieved through graph message-passing encoders, where geometric features (i.e., the “message”) are exchanged only between directly connected nodes within the same connected component of a graph—defined in this case by the polygon’s linear rings.
+As a result, PolyMP-based models can better preserve structural relationships in polygons with complex topologies.
 6.2 Limitations
-Of course, PolyMP still encounters limitations in terms of misclassification.
-Fig.
-8.
-demonstrates failures of PolyMP classifications on the Glyph and OSM datasets.
-On synthetic data, misclassifications occur predominantly on samples where excessive geometric transformations (e.g., heavy shearing and anisotropic scaling) destroy the gestalt of the shape to the extent it is no longer label preserving (Fig.
-8(a): letter V, Z and R).
-A special case are shapes with different labels dependent on orientation – M and W, where depending on the extent of rotation and shearing, misclassifications occur.
-For instance in row two of Fig.
-8(a), we observe that rotated letters W can be misclassified into letter M, or even Z; heavily sheared W can be misclassified as H and V.
-Similarly for building polygons (OSM), PolyMP may misclassify polygons sharing similar geometric shapes (E, vs L or U) (Fig.
-8).
+Despite its strong performance, our proposed method still encounters misclassification limitations.
+Figure 9 highlights instances where PolyMP fails to classify samples correctly in the Glyph and OSM datasets.
+On synthetic data, misclassifications predominantly occur in cases where excessive geometric transformations—such as heavy shearing and anisotropic scaling—severely distort the shape’s structure, rendering it unrecognizable: rotated W shapes (Fig.9a) are sometimes mistaken for M or even Z, while heavily sheared W shapes may be misclassified as H or V.
+Similarly, in the OSM dataset, PolyMP and PolyMP-DSC occasionally misclassify building polygons with similar geometric structures, such as E, L, and U shapes (Fig.
+9b).
+This can be attributed to the model’s reliance on local geometric features, which may not sufficiently distinguish between structures with high shape similarity.
 Considering the characteristics of the two datasets (Sec.
-3.1), PolyMP trained Glyph dataset learns locally salient geometric features via message-passing layers from serif glyphs, which contain salient decorative strokes ending.
-Letters sharing similar geometric shapes are misclassified when the geometric shapes of classified polygons are destroyed by excessive geometric transformations.
+4.1 and 4.2), PolyMP and PolyMP-DSC trained on the Glyph dataset primarily learns locally salient geometric features through message-passing encoders.
+These features are particularly effective for serif glyphs, which contain distinctive decorative stroke endings.
+However, when excessive geometric transformations distort these key features, the model struggles to preserve label consistency, leading to misclassification among shapes with similar structures.
 7 Conclusion
-Based on extensive empirical evidence given by the analysis of invariances to geometric transformations, trivial vertices and model generalizability, we demonstrate that discrete, non-grid like data representations in combination with graph-based message-passing neural networks (PolyMP) are optimal for deep learning models on 2D vector geometries.
-We demonstrate a robust ability to learn shapes of polygonal geometries, and demonstrate the ability to generalize to a real-world building footprint dataset.
-We find that a learning model based on graph data representation that explicitly encodes both node and explicit edge structures leads to a robust approach enabling highly accurate feature learning on geometric shapes – conveniently, these structures are readily available from the 2D simple feature representations of geometries implemented in current GIS libraries.
-Our results indicate that representations play an important part providing a strong inference bias for geometric shape learning, and tasks such as the classification of polygons and boosting model performance.
-We believe that our approach to learning geometric-invariant features of polygons is not limited to simple geometric shape classification.
-This approach can benefit suitable geospatial tasks where the geometric-invariant features of polygons matches the learning objectives (i.e., geospatial regression tasks involving spatial relation predictions of polygon map [8]).
-We also discuss the limitations of PolyMP through misclassification examples on datasets.
-Through the failure cases of classification, we observe that the message-passing network for graph data representation of polygons has limited capability of transferring locally salient geometric features from one domain to another.
-The simple message-passing network has limited control on learning/not learning geometric invariant features as need.
-The geometric invariant feature learning of vector geometries is currently an area of intense investigation in spatial deep learning [2, 1, 6, 35, 8].
-Unlike in learning based on raster datasets or point clouds, the segmentation problem is partially solved by the explicit encoding of structured vector geometries.
-Here, we show results that allow us to undertake the classification of spatial geometric shapes efficiently, based on decisions that can be grounded in theory and are experimentally validated.
-In future work, we will further examine how richer graph encoding of the structure of geometries, as well as of the relationships between proximal geometries could further boost model performance, or enable the segmentation of complex spatial object arrangements.
-Statements and Declarations
-Data availability. The data and codes that support the findings of this study are available with the identifier(s) at the private link: https://figshare.com/s/b781d7724d3220e24c7c.
-Conflicts of interest. This paper has been approved by all co-authors. The authors have no competing interests to declare that are relevant to the content of this article.
+In this study, we investigated the challenge of geometric-invariant shape classification from 2D vector polygons, a common but underexplored data format in spatial analysis.
+Through comprehensive empirical evaluation, we demonstrated that combining discrete, non-grid-like polygon representations with graph-based message-passing neural networks—PolyMP and its densely self-connected variant, PolyMP-DSC—enables the learning of robust and expressive geometric features.
+Our findings show that graph representations of polygons, which encode both geometric structure and vertex connectivity, significantly outperform sequence- and set-based encodings in terms of transformation invariance and generalization.
+In particular, PolyMP-DSC enhances feature propagation by incorporating dense self-connections, resulting in improved classification performance and robustness to structural perturbations such as trivial vertices.
+These properties are crucial for geospatial tasks where object shapes may undergo rotation, scaling, shearing, or digitization noise.
+Importantly, our approach aligns well with practical geospatial workflows: polygon-to-graph conversion is natively supported by many GIS libraries, enabling seamless adoption of our models in downstream applications such as automated cartographic generalization, building footprint recognition, and road geometry analysis.
+Nevertheless, current limitations remain.
+Both PolyMP and PolyMP-DSC occasionally struggle with locally distinctive features that vary across shape classes or domains.
+To address this, future work will explore the integration of complementary geometric-invariant descriptors—such as spectral features (e.g., NUFT) and local shape attributes (e.g., turning angles, curvature, or radii of arc segments)—to further improve model robustness and invariance.
+Overall, this work highlights the importance of representation choice in spatial deep learning.
+By designing models with inductive biases that reflect geometric structure and transformation invariance, we move closer to human-like perception of shape in geospatial domains.
+Beyond classification, our findings open promising directions for extending graph-based learning to polygon regression, spatial clustering, and topological inference in large-scale vector datasets.
+Declarations
+Data availability.
+The data and code supporting the findings of this study are available at https://github.com/zexhuang/PolyMP.
+Conflicts of interest.
+This paper has been approved by all co-authors. The authors have no competing interests to declare that are relevant to the content of this article.
+Funding.
+No funding was obtained for this study.